### v2

# PDE-Transformer: A Continuous Dynamical Systems Approach to Sequence Modeling

Abstract
We propose PDE-Transformer, a novel sequence‐modeling paradigm that casts the forward pass of a Transformer as the numerical discretization of a continuous reaction–diffusion system derived from a variational energy functional. In our framework, token embeddings evolve under a partial differential equation whose nonlocal integral term models self-attention, local reaction term models feed-forward layers, diffusion term encodes positional smoothing, and a stability control term corresponds to layer normalization. From this unifying perspective, we design an Adaptive PDE Diffusion Layer—an efficient, learnable finite-difference stencil that enforces local smoothness in feature space with linear time complexity and complements self-attention’s global routing. Through a systematic theoretical analysis based on four pillars (stability, diffusion geometry, multi-scale dynamics, and component coupling), we derive principled guidelines for integrating the PDE layer at seven candidate points in the Transformer. Empirically, on the Long Range Arena benchmark, placing the layer immediately after embedding yields a 4.1 pp average accuracy gain over a strong baseline, and an adaptive multi-scale variant delivers further improvements. Our work thus offers a principled, lightweight mechanism to bolster long-range dependency modeling by harmonizing continuous PDE smoothing with discrete self-attention.
1 Introduction
Since its inception, the Transformer architecture has achieved revolutionary breakthroughs across diverse fields, from natural language processing and computer vision to protein folding, owing to its powerful self-attention mechanism. However, as applications scale from short texts to ultra-long sequences—such as document-level dialogues or whole-genome sequences—the standard Transformer reveals two fundamental bottlenecks. First, its computational and memory complexity, which scales quadratically with sequence length $L$ ($O(L^{2})$), becomes prohibitively expensive. Second, its purely content-driven global interaction mechanism lacks explicit modeling of local geometric structure, hindering its ability to effectively capture long-range dependencies. Although variants like sparse attention and low-rank approximations offer patchwork optimizations, they remain engineering modifications to a discrete computational graph, failing to transcend the limitations of the discrete paradigm or provide a unified theoretical framework for their design.
To overcome this impasse, we propose a return to the first principles of physics, reframing the process of sequence modeling as a continuous variational dynamical system. In this framework, token embeddings do not undergo discrete layered transformations but rather evolve within an energy field governed by three fundamental forces: local diffusion, nonlinear reaction, and nonlocal coupling. We establish a one-to-one correspondence between these forces and the core components of the Transformer: the nonlinear reaction maps to the feed-forward network (FFN), and the nonlocal coupling maps to the self-attention mechanism. This perspective strikingly reveals a structural deficiency in the standard Transformer: the absence of the crucial diffusion term. In physical systems, this term is responsible for penalizing sharp variations and imposing local smoothness, which is key to forming stable, ordered structures.
Building on this core finding, we design a lightweight, plug-and-play Adaptive PDE Diffusion Layer with linear complexity. This module discretizes a reaction-diffusion equation via a learnable finite-difference method, introducing a structured local smoothness inductive bias into the model. It forms a natural complementarity with self-attention: the PDE diffusion layer is responsible for reinforcing local geometric consistency, while self-attention focuses on capturing global content-based associations.
To validate this paradigm, we systematically investigate the impact of integrating the PDE diffusion layer at seven distinct points within the Transformer architecture. Experiments on the challenging Long Range Arena (LRA) benchmark provide compelling evidence for our theory: placing the PDE layer immediately after the initial token embeddings and before the first Transformer block yields the most significant performance gains. This configuration achieves an average accuracy improvement of 4.1 percentage points over a strong baseline, with a multi-scale variant delivering further improvements. This finding uncovers a core mechanism: before global, sparse attention interactions can be effective, the raw semantic space must first undergo local, dense, structured smoothing.
The main contributions of this work are fourfold. First, we introduce a new theoretical paradigm that unifies sequence modeling as a continuous reaction–diffusion dynamical system derived from variational principles, offering a novel physical lens to understand and improve Transformer architectures. Second, we design the Adaptive PDE Diffusion Layer as a plug-and-play, linear-complexity module that significantly enhances the model’s ability to capture local structure at negligible computational cost. Third, we conduct a systematic empirical study to identify the optimal integration point for such a local smoothing mechanism within the Transformer. Finally, we reveal the profound complementarity between explicit local geometric smoothing and global content-based aggregation, providing principled guidance for building more robust and efficient long-sequence models.
The remainder of this paper is organized as follows. Section 2 reviews prior work on efficient Transformers and continuous-time models. Section 3 details our theoretical framework, deriving the PDE-Transformer from first principles. Section 4 presents our comprehensive experimental validation on the LRA benchmark. Finally, Section 5 concludes the paper and discusses future directions.
2 Related Work
We situate the PDE-Transformer within three major research streams: (i) efficiency enhancements to discrete Transformers, (ii) continuous-time sequence modeling, and (iii) neural networks as PDE solvers. Together, these perspectives highlight both the progress and the remaining gaps that motivate our work.
2.1 Discrete Efficiency Enhancements
The quadratic complexity of the vanilla Transformer Vaswani et al. (2017) has motivated extensive efforts to improve scalability. One class of methods reduces attention cost by introducing fixed or learned sparsity, such as local windows combined with global tokens Beltagy et al. (2020); Zaheer et al. (2020) or strided/dilated patterns Child et al. (2019). While effective, these approaches impose pre-defined structures that may not align with data and risk bottlenecks in capturing long-range dependencies. A second class leverages low-rank or kernel approximations: Linformer Wang et al. (2020) projects Key and Value matrices into lower-dimensional spaces, while Performer Choromanski et al. (2021) approximates softmax attention with kernel features. These methods reduce complexity but rely on restrictive low-rank assumptions that can discard fine-grained information in the long-tail spectrum. A third family introduces recurrence and memory, such as Transformer-XL Dai et al. (2019), which caches hidden states across segments to extend context length. This alleviates quadratic cost but introduces challenges in compressing historical states and handling discontinuities at segment boundaries. Overall, these methods optimize the discrete computation graph—deciding “who attends to whom”—without altering the fundamental discrete paradigm (see Table 1).
Table 1: Taxonomy of efficient Transformer architectures. These methods improve scalability within a discrete framework but do not change the underlying paradigm.
Approach Key Examples Core Idea Limitation
Sparsity BigBird Zaheer et al. (2020), Longformer Beltagy et al. (2020) Restrict attention to sparse patterns Rigid information flow; may miss dependencies
Low-Rank / Kernel Linformer Wang et al. (2020), Performer Choromanski et al. (2021) Approximate the full attention matrix Low-rank assumption; loss of high-frequency info
Recurrence / Memory Transformer-XL Dai et al. (2019) Segment sequence and reuse past states State compression bottleneck; boundary effects
2.2 Continuous-Time Models
A more fundamental line of research interprets neural networks as discretizations of continuous-time systems. Neural ODEs Chen et al. (2018) recast residual networks as solutions of ODEs, treating depth as continuous and enabling benefits such as memory-efficient training Gholami et al. (2019) and irregular time-series modeling Li et al. (2020). However, these methods continuous-ize the depth dimension rather than the sequence.
State-Space Models (SSMs) directly target the sequence dimension. They model dynamics with a linear ODE:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t)+\mathbf{D}x(t),$ (1)
where $x(t)$ is the input, $h(t)$ the latent state, and $y(t)$ the output. The S4 model Gu et al. (2022) demonstrated that by structuring $\mathbf{A}$, one can discretize this system into an efficient convolutional form. Recent advances like Mamba Gu & Dao (2023) incorporate selection mechanisms and input-dependent parameters, surpassing linear time-invariant constraints and achieving strong performance. SSMs thus represent the most competitive continuous-time alternative to Transformers. However, they are governed by ODEs, which describe the temporal evolution of state vectors. Our PDE formulation instead models a field $u(x,t)$, introducing explicit spatial derivative terms (e.g., $\nabla^{2}u$) and enabling spatiotemporal coupling along the sequence dimension (Table 2).
Table 2: Comparison of continuous-time modeling paradigms. Unlike ODE-based models, our PDE approach introduces explicit spatiotemporal coupling.
Paradigm Continuous Dim. Equation Mechanism Limitation
Neural ODE Chen et al. (2018) Depth ODE Continuous layers Models depth, not sequence
SSM (Mamba) Gu & Dao (2023) Sequence ODE Temporal evolution + selection Lacks explicit spatial coupling
PDE-Transformer (Ours) Sequence PDE Spatiotemporal coupling Models direct interaction via $\nabla^{2}u$
2.3 Neural Networks as PDE Solvers
A parallel literature connects neural networks and PDEs but with inverse objectives. Physics-Informed Neural Networks (PINNs) Raissi et al. (2019) parameterize PDE solutions with neural networks, enforcing boundary conditions and PDE residuals during training. Neural Operators such as the Fourier Neural Operator (FNO) Li et al. (2021) extend this idea by learning resolution-invariant mappings from input functions to PDE solutions, effectively serving as universal solvers. In contrast, our work does not seek to solve externally-given PDEs. Instead, we adopt PDE principles—diffusion, waves, and reaction—as internal design guidelines for the information dynamics of sequence models.
2.4 Synthesis
In summary, discrete efficiency methods enhance scalability without altering the static computation graph; continuous-time models, particularly SSMs, provide strong alternatives but are ODE-based and lack explicit spatial coupling; and neural PDE solvers apply PDEs as external constraints rather than intrinsic dynamics. This landscape reveals an open frontier: applying PDEs as first-principles foundations for neural sequence architectures. Our PDE-Transformer advances this direction by formalizing spatiotemporal coupling as a core design principle for long-range modeling.
3 Theory and Method: Unified Variational Dynamical System and Adaptive Diffusion
To address the dual bottlenecks of quadratic complexity and the lack of local geometric modeling in Transformers, we reframe sequence modeling as a variational dynamical system from first principles. In this view, representations evolve under local diffusion, nonlinear reaction, and nonlocal coupling forces. This reveals a structural deficiency in standard Transformers—the missing diffusion term—which we remedy with a practical and efficient adaptive PDE diffusion layer. In this section, we derive the governing dynamics from an energy functional, map its components to Transformer modules, and instantiate the theory as a trainable layer, further proposing a principled integration strategy and discussing its multiscale spectral properties.
3.1 Unified Variational Principle
(a) Unified variational framework: energy functional decomposition and gradient-flow PDE.
(b) Architecture mapping: PDE components vs. standard Transformer, showing the missing diffusion term and our proposed PDE diffusion layer.
Figure 1: Theory framework of PDE-Transformer. Two stacked panels show (a) the unified variational formulation and corresponding PDE, and (b) the architectural mapping that highlights the missing diffusion component in standard Transformers.
We denote the continuous representation of the sequence as $u(x,t)$, where $x\in\Omega$ represents the position and $t$ the “evolution time” corresponding to network depth. The global energy functional is defined as
$E[u]$ $=\int_{\Omega}\Big(\tfrac{\alpha}{2}|\nabla u|^{2}+F(u)\Big)dx+\frac{\beta}{2}\iint_{\Omega\times\Omega}K(x,y)[u(x)-u(y)]^{2}dxdy.$ (2)
It consists of three parts: a local tension term $|\nabla u|^{2}$ that penalizes sharp variations to maintain local consistency; a nonlinear potential $F(u)$ encoding task objectives; and a long-range coupling $K(x,y)$ establishing content-dependent global interactions.
Theorem 3.1 (Unified Dynamical Equation).
The gradient flow of $E[u]$ is
$\frac{\partial u}{\partial t}=\alpha\Delta u-F^{\prime}(u)-\beta\mathcal{L}_{K}[u],\qquad\mathcal{L}_{K}[u](x)=\int K(x,y)[u(x)-u(y)]dy.$ (3)
Each term corresponds directly to a Transformer module. The standard Transformer realizes reaction and nonlocal coupling but ignores diffusion, limiting explicit geometric modeling and stability.
Table 3: Physical and functional correspondence of PDE-Transformer components.
Mathematical Term Physical Process Module Function Information Pattern
$\alpha\Delta u$ Diffusion PDE Diffusion Layer (Ours) Regularizes locally Geometric, dense
$F^{\prime}(u)$ Reaction Feedforward Network Pointwise transformation Local, independent
$\beta\mathcal{L}_{K}[u]$ Nonlocal coupling Self-Attention Aggregates globally Sparse, content-driven
3.2 Diffusion Dynamics and Stability
To align theory with implementation, we adopt Neumann (reflective) boundary conditions:
$(\mathbf{\Delta}_{N}X)_{i}=\begin{cases}X_{2}-X_{1},&i=1,\\ X_{i-1}-2X_{i}+X_{i+1},&2\leq i\leq L-1,\\ X_{L-1}-X_{L},&i=L.\end{cases}$ (4)
This preserves boundary information and is compatible with discrete cosine transform (DCT) analysis.
Theorem 3.2 (Spectrum).
The eigenvalues of $\mathbf{\Delta}_{N}$ are $\lambda_{k}=-4\sin^{2}(\pi k/(2L))$ with cosine eigenvectors.
Corollary 3.3 (CFL Condition).
For the explicit iteration $X^{(k+1)}=X^{(k)}+\alpha\mathbf{\Delta}_{N}X^{(k)}$, stability requires $\alpha<\tfrac{1}{2}$.
Theorem 3.4 (Lyapunov Monotonicity).
Under the CFL condition, the Dirichlet energy $V(X)=X^{\top}(-\mathbf{\Delta}_{N})X$ decreases monotonically, ensuring stability and improving the optimization landscape.
3.3 Information-Theoretic Integration
Consider a forward chain $X_{0}\to X_{1}\to\cdots\to X_{N}\to Y$. For any smoothing operator $\mathcal{S}$, define the information retention:
$\rho(d)=\mathbb{E}\left[\frac{I(\mathcal{S}(X^{(d)});Y)}{I(X^{(d)};Y)}\right],\qquad\rho(1)\geq\rho(2)\geq\cdots\geq\rho(N).$ (5)
This shows that earlier diffusion preserves more task-relevant information.
We further compare seven candidate insertion points using the value function $V_{i}=w_{I}I_{i}-w_{D}D_{i}-w_{C}C_{i}$, where $I_{i}$ denotes information retention, $D_{i}$ distortion, and $C_{i}$ computational cost.
Table 4: Comparison of candidate insertion points for diffusion layers.
Pos. Mechanism Property Info Dist. Cost Rank
1 After Embedding $\|X_{0}^{(s)}-X_{0}\|=O(\alpha)$ High Low Low 1
2 After MLP Local proximal step High Low Low 2
3 Between Layers Gradient flow improvement Mid-High Mid Mid 3
4 Before LN Non-commutativity Mid Mid Low 4
5 Inside Attention Breaks normalization Mid-Low Mid-High Mid 5
6 Between Heads Loss of diversity Low High Mid 6
7 After Attention Breaks sparsity Low High Low 7
Overall Ranking: After embedding >> After MLP >> Between layers >> Before LN >> Inside attention >> Between heads >> After attention.
3.4 Multi-Scale Spectral Complementarity
Figure 2: Frequency domain analysis of the multi-scale diffusion mechanism. (Left) The transfer function $H(\omega)$ for different diffusion scales. Single-scale diffusions (Fast, Medium, Slow) act as low-pass filters with different cutoff frequencies. (Right) The energy distribution across four frequency bands. The multi-scale approach (green, dashed line) achieves a more balanced energy distribution across the entire frequency spectrum compared to any single-scale method, enabling it to capture a richer set of signal components from both global trends (low frequency) and local details (high frequency).
For diffusion with step size $h$, the frequency response is
$H_{h}(\omega)=1-4\alpha\sin^{2}\!\Big(\tfrac{\omega h}{2}\Big).$ (6)
Small steps ($h=1$) preserve high-frequency details, while larger steps ($h=2,4$) emphasize mid-to-low frequencies. Mixing scales covers the entire spectrum; a practical initialization uses weights $1:0.6:0.3$. Diffusion is geometry-driven and dense, with effective radius proportional to $\sqrt{t}$, while attention is content-driven and sparse, connecting arbitrary positions. Their combination captures both local geometry and global dependency.
3.5 Adaptive PDE Diffusion Layer
The proposed module incorporates multi-scale diffusion ($[1,2,4]$), learnable coefficients per scale and channel, and CFL-enforced constraints $\alpha<0.5$ at runtime. Its complexity is $O(Ld)$ for a single scale or $O(KLd)$ for multiple scales, significantly lower than attention $O(L^{2}d)$. Theory and information analysis indicate that the best insertion point is after the embedding layer, before the first Transformer block, maximizing information retention and stability while improving optimization.
Summary. We unify sequence modeling into a reaction–diffusion–nonlocal coupling system, derive its integro-PDE gradient flow, and identify the missing diffusion term in Transformers. Spectral and stability analysis motivates an efficient adaptive PDE diffusion layer, with its optimal position after embedding. This principled module augments global coupling with explicit local modeling, laying the foundation for subsequent experiments. applications.
4 Experiments
This section aims to empirically validate the core claims of our theoretical framework through a series of rigorous analyses. Our experimental design follows a clear logical progression: first, we systematically compare different integration strategies to pinpoint the optimal mechanism for the Adaptive PDE Diffusion Layer, thereby adjudicating the theoretical debate from Section 4. Second, we conduct an ablation study to verify the effectiveness of the multi-scale dynamics. All experiments are conducted on a challenging long-sequence benchmark to provide a stringent test of our theory.
4.1 Experimental Setup
We conduct all experiments on the Long Range Arena (LRA) benchmark Tay et al. (2020), which includes five tasks spanning diverse modalities and challenges. Our experimental models integrate the proposed Adaptive PDE Diffusion Layer into a strong, optimized vanilla Transformer baseline. We systematically evaluate seven distinct integration positions, as detailed in Table 5, to determine the optimal placement. All models were trained on NVIDIA A100-80GB GPUs, with reproducibility ensured by fixing random seeds. Comprehensive details regarding the datasets, task-specific model configurations, unified training hyperparameters, and the software/hardware environment are provided in Appendix A to ensure full transparency and reproducibility of our work.
Table 5: Design space of PDE integration positions in Transformer architectures.
Pos Integration Point Description
1 After Embedding Applied right after the input embedding layer, before any Transformer block, so diffusion acts on raw semantic representations.
2 After MLP Inserted after each block’s MLP sub-layer; diffusion runs on features that have passed through attention and feed-forward transformations.
3 Layer Diffusion Placed between consecutive Transformer layers to promote inter-layer information flow across depth.
4 Before LayerNorm Applied just before LayerNorm in every sub-layer, operating on unnormalized features.
5 In Attention Integrated into the attention mechanism itself—e.g., on attention weights or intermediate values—during computation.
6 Head Diffusion Acts across attention heads within the same layer, enabling information exchange among heads.
7 After Attention Inserted immediately after the self-attention sub-layer and before the MLP, diffusing attention-processed representations.
4.2 Main Results: Pinpointing the Optimal PDE Mechanism
Table 6: Average accuracy on the LRA benchmark for different PDE integration positions. Detailed per-task scores are in Appendix B.
Integration Position Avg. Accuracy
PDE-After-Embedding 0.6269
PDE-After-MLP 0.5986
PDE-Layer-Diffusion 0.5970
PDE-Before-LayerNorm 0.5962
PDE-In-Attention 0.5909
PDE-Head-Diffusion 0.5884
Baseline Transformer 0.5862
PDE-After-Attention 0.5617
To systematically adjudicate the theoretical debate from Section 4 regarding the optimal integration position, we conducted a comprehensive comparison of seven different PDE configurations.Table 6 summarizes the average performance of each configuration on the LRA benchmark.
The results clearly show that the ‘After Embedding‘ position achieves a compellingly superior performance, outperforming the strong baseline by 4.07 percentage points. This finding provides powerful empirical support for our proactive pre-processing hypothesis. It reveals a core mechanism: applying Pillar II (Diffusion Geometry) directly on the initial, potentially noisy semantic manifold enforces a local smoothness inductive bias that provides a more regularized and robust foundation for all subsequent non-local interactions (i.e., self-attention). This strategy of “semantic regularization” at the source is substantially more effective than “reactive refinement” later in the information flow (e.g., ‘After MLP‘).
Concurrently, the failure of the ‘After Attention‘ position is equally insightful. It delineates the boundaries of our method: once a component (self-attention) has learned valuable sparse association patterns, forcibly imposing dense local smoothing can act as a destructive interference. This observation reinforces the functional complementarity between PDE diffusion and self-attention discussed in Section 4.1—they are orthogonal mechanisms, and their simple serial composition is not always optimal, pointing towards future explorations of more sophisticated fusion strategies as suggested by Pillar IV.
Figure 3: Overall performance comparison on the LRA benchmark. Error bars show standard deviation across five runs (n=5n{=}5).
4.3 Ablation Study: Validating Multi-Scale Dynamics
Having identified ‘After Embedding‘ as a superior integration strategy, we conducted a further ablation study to validate the efficacy of Pillar III (Multi-Scale Dynamics). We selected the top three performing positions and compared the effects of different diffusion scales ($h=1,2,4$) versus the adaptive multi-scale combination on the ListOps task. The results are presented in Table 7.
Table 7: Ablation study results for the multi-scale diffusion mechanism on the ListOps task.
Position Fast Medium Slow Multi-
(hh=1) (hh=2) (hh=4) scale
After Emb. 0.3960 0.3940 0.3990 0.4080
After MLP 0.3900 0.3930 0.3910 0.4010
Layer Diff. 0.3850 0.3890 0.3910 0.3970
The results are remarkably consistent: for all tested positions, the adaptive ‘Multi-scale‘ configuration outperforms any single scale. This strongly corroborates our theoretical hypothesis that real-world sequential data contains structural information at varying granularities. By enabling the model to adaptively combine different diffusion processes—capturing high-frequency local details via short-range diffusion and establishing low-frequency global trends via long-range diffusion—it can learn more comprehensive and robust feature representations. This empirical success is also in strong agreement with frequency analysis in signal processing theory (see Appendix C.3).
Figure 4: Detailed results of the multi-scale PDE ablation study on the ListOps task, broken down by PDE position.
4.4 Conclusion of Experiments
In summary, our empirical investigation systematically validates the core tenets of our theoretical framework, culminating in a clear and powerful conclusion: **introducing a local smoothness inductive bias to Transformers is an effective design principle, and the optimal strategy to achieve this is through proactive pre-processing of input representations with an adaptive, multi-scale PDE diffusion layer.** This conclusion not only demonstrates the effectiveness of the proposed PDE-Transformer but, more importantly, it illuminates the profound complementary relationship between the global information aggregation of self-attention and the local structural smoothing of PDE diffusion. Our work offers a new, principled, and rigorously validated approach for future sequence model design: ensure a robust understanding of underlying local structure before modeling complex long-range dependencies.
5 Conclusion
PDE-Transformer introduces a fundamentally new way to think about sequence modeling by reframing the Transformer’s forward pass as the discretization of a continuous reaction–diffusion system; by deriving an energy functional whose gradient flow yields a PDE with four natural components—non-local interaction for self-attention, local reaction for feed-forward networks, diffusion for positional smoothing, and stability control for normalization—it provides a unified theoretical lens that explains why residual connections and layer normalization are not mere engineering tricks but necessary mechanisms for well-posedness and stability, bridging gaps in our understanding of long-range dependency modeling and revealing that diffusion processes yield polynomial-decay kernels that more effectively capture distant interactions than the exponential decay inherent in standard attention; building on this theory, an Adaptive PDE Diffusion Layer is designed that approximates the Laplacian via second-order finite differences with a learnable diffusion coefficient to adaptively enforce local smoothness, and when integrated into the Transformer at the optimal location—immediately after the embedding layer—this lightweight module achieves a 4.1 pp average accuracy boost on the Long Range Arena benchmark, with a multi-scale version delivering further gains, while a thorough theoretical analysis of seven integration strategies and extensive experiments demonstrate that continuous PDE smoothing and global self-attention are highly complementary, offering a principled, efficient route to robust, long-sequence modeling.
6 Ethics Statement
Our research does not involve human subjects, personally identifiable information, or sensitive data. All datasets used in experiments are publicly available and have been widely adopted in prior work. We are not aware of any foreseeable misuse or harmful applications directly stemming from our proposed methods. We further confirm that there are no conflicts of interest, sponsorship concerns, or legal compliance issues associated with this work. We acknowledge that any potential societal impact of large-scale models, such as fairness, bias, or privacy concerns, lies beyond the specific scope of this paper but remains an important consideration for future research.
7 Reproducibility Statement
We place strong emphasis on the reproducibility of our results. In the main text, we clearly describe our proposed theoretical framework, its derivation, and the algorithmic implementation details. A complete set of proofs for theoretical results is provided in the Appendix. Hyperparameter configurations, training setups, and ablation study details are explicitly reported in the Experiments section. All datasets used are standard and publicly available, and their preprocessing steps are fully documented in the supplementary material. To facilitate replication, we will release anonymized source code and scripts for reproducing all experiments upon publication.
8 LLM Usage Statement
In accordance with ICLR guidelines on the disclosure of Large Language Model (LLM) usage, we clarify that no LLM contributed substantively to the conception, methodology, or analysis presented in this paper. LLMs (e.g., ChatGPT) were used exclusively as auxiliary tools for writing assistance, language refinement, and stylistic editing. All technical content, theoretical contributions, experimental design, and analysis were conceived, implemented, and validated entirely by the authors. The role of LLMs was limited to improving clarity of presentation and does not rise to the level of authorship or contribution under ICLR policy.
Appendix A Appendix: Rigorous Mathematical Proofs
This appendix provides the rigorous mathematical proofs for the theoretical frameworks underpinning the PDE-Transformer. These proofs establish the model’s stability, long-range modeling capability, multi-scale representation power, and component coordination.
A.1 A. Theoretical Framework I: Lyapunov Stability
This framework analyzes the stability of the model’s underlying gradient flow system using Lyapunov’s second method.
Theorem A.1 (Energy Monotonicity and Global Stability).
Consider the energy functional $E[u]$ within the PDE-Transformer, defined as:
$E[u]=\int_{\Omega}\left(\frac{1}{2}|\nabla u|^{2}+F(u)+\frac{\lambda}{2}|u-u_{0}|^{2}\right)dx,$
which corresponds to the gradient flow dynamics:
$\frac{\partial u}{\partial t}=-\frac{\delta E}{\delta u}=\Delta u-F^{\prime}(u)-\lambda(u-u_{0}).$
If the potential $F(u)$ is strictly convex, satisfying $F^{\prime\prime}(u)\geq\mu>0$, the system is globally asymptotically stable.
Proof.
We establish stability by demonstrating that the energy functional $E[u]$ serves as a strict Lyapunov function for the system.
Step 1: Construct the Lyapunov Function. We select the Lyapunov function $V[u]=E[u]$. By choosing an appropriate reference, we can assume $F(u)\geq 0$. Since $\lambda>0$, the functional $V[u]$ is bounded from below by 0.
Step 2: Compute the Time Derivative. We compute the time derivative of $V[u]$ along the system’s trajectories:
$\frac{dV}{dt}$ $=\int_{\Omega}\frac{\delta E}{\delta u}\frac{\partial u}{\partial t}\,dx$
$=\int_{\Omega}\frac{\delta E}{\delta u}\left(-\frac{\delta E}{\delta u}\right)\,dx$
$=-\int_{\Omega}\left\|\frac{\delta E}{\delta u}\right\|^{2}\,dx\leq 0.$
Equality holds if and only if $\frac{\delta E}{\delta u}=0$, which defines the equilibrium state of the system.
Step 3: Apply LaSalle’s Invariance Principle. The condition $F^{\prime\prime}(u)\geq\mu>0$ ensures that $E[u]$ is strictly convex, guaranteeing a unique global minimum. Since $V[u]$ is bounded below and its time derivative is negative semi-definite, all trajectories must converge to the largest invariant set where $\frac{dV}{dt}=0$. This set consists solely of the unique global minimum. Thus, the system is globally asymptotically stable. ∎
Theorem A.2 (Exponential Decay of the Gradient Norm).
Under the conditions of Theorem A.1, the $L^{2}$-norm of the gradient decays at an exponential rate:
$\left\|\frac{\delta E}{\delta u}(t)\right\|_{L^{2}}\leq\left\|\frac{\delta E}{\delta u}(0)\right\|_{L^{2}}e^{-\mu t}.$
Proof.
Step 1: Define the Evolution Equation for the Gradient. Let $v=\frac{\partial u}{\partial t}=-\frac{\delta E}{\delta u}$. Taking the time derivative of $v$ yields:
$\frac{\partial v}{\partial t}=\frac{\partial}{\partial t}\left(-\frac{\delta E}{\delta u}\right)=\Delta v-F^{\prime\prime}(u)v-\lambda v.$
Step 2: Analyze the Dynamics of the Gradient Norm. We examine the time derivative of the squared $L^{2}$-norm of $v$:
$\frac{1}{2}\frac{d}{dt}\|v\|_{L^{2}}^{2}$ $=\int_{\Omega}v\frac{\partial v}{\partial t}\,dx$
$=\int_{\Omega}v(\Delta v-F^{\prime\prime}(u)v-\lambda v)\,dx$
$=-\|\nabla v\|_{L^{2}}^{2}-\int_{\Omega}F^{\prime\prime}(u)v^{2}\,dx-\lambda\|v\|_{L^{2}}^{2}.$
Step 3: Establish the Differential Inequality. Using the condition $F^{\prime\prime}(u)\geq\mu$ and dropping the non-positive term $\|\nabla v\|_{L^{2}}^{2}$, we obtain:
$\frac{d}{dt}\|v\|_{L^{2}}^{2}\leq-2\mu\|v\|_{L^{2}}^{2}.$
Applying Grönwall’s inequality to this differential inequality yields $\|v(t)\|_{L^{2}}^{2}\leq\|v(0)\|_{L^{2}}^{2}e^{-2\mu t}$. Taking the square root of both sides completes the proof. ∎
A.2 B. Framework II: Diffusion Geometry
Theorem A.3 (Polynomial Decay of the Heat Kernel).
The heat kernel $K_{t}(x,y)$ corresponding to the 1D discrete Laplacian operator exhibits an asymptotic behavior described by a Gaussian:
$|K_{t}(x,y)|\leq Ct^{-1/2}\exp\left(-\frac{|x-y|^{2}}{4Dt}\right).$
This implies a fundamentally slower, polynomial-like decay of influence over long distances compared to the exponential decay of standard self-attention mechanisms.
Proof.
Step 1: Spectral Representation. The proof relies on the spectral representation of the heat kernel:
$K_{t}(x,y)=\sum_{k}e^{-\lambda_{k}t}\phi_{k}(x)\phi_{k}(y),$
where $\lambda_{k}$ and $\phi_{k}$ are the eigenvalues and eigenfunctions of the discrete Laplacian.
Step 2: Asymptotic Analysis. For large time $t$ or small grid spacing, the discrete sum can be approximated by an integral via the Poisson summation formula. This integral evaluates to the continuous heat kernel, which is a Gaussian function. While the Gaussian itself has an exponential tail, its integrated effect over time results in a polynomial decay of influence, which is fundamentally slower than the direct exponential decay $e^{-c|x-y|}$ found in typical attention mechanisms. ∎
A.3 C. Framework III: Multi-Scale Dynamics
Theorem A.4 (Multi-Scale Approximation Error Bound).
Let $\Delta_{\text{multi}}=\sum_{k=1}^{K}\alpha_{k}\Delta_{h_{k}}$ be a multi-scale approximation of the ideal Laplacian $\Delta$. By selecting scales $\{h_{k}\}$ geometrically and optimizing weights $\{\alpha_{k}\}$, the approximation error is bounded by:
$\left\|\Delta_{\text{multi}}-\Delta\right\|_{\mathcal{L}(H^{s},H^{s-2})}\leq CK^{-p},$
where $p\geq 1$ is the order of convergence.
Proof.
This theorem is a result from numerical analysis and approximation theory. The problem can be framed as approximating the function $f(\omega)=-\omega^{2}$ (the symbol of the ideal Laplacian) with a linear combination of basis functions $\hat{\Delta}_{h_{k}}(\omega)=-\frac{4}{h_{k}^{2}}\sin^{2}(\frac{\omega h_{k}}{2})$. By choosing scales $h_{k}$ in a geometric progression, the basis functions effectively tile the frequency domain. Techniques from Chebyshev approximation or spline theory can then be used to show that the $L^{2}$ error of the best approximation decreases polynomially with the number of scales $K$. ∎
A.4 D. Framework IV: Multi-Component Coupling
Theorem A.5 (Synchronization of Coupled Systems).
Consider a system of $H$ coupled PDEs for multiple attention heads:
$\frac{\partial u_{i}}{\partial t}=\alpha_{i}\Delta u_{i}+\sum_{j\neq i}\beta_{ij}(u_{j}-u_{i})+f_{i}(u_{i}).$
If the coupling graph defined by $\beta_{ij}>0$ is connected, the system asymptotically synchronizes, i.e., $\lim_{t\to\infty}\|u_{i}(t)-u_{j}(t)\|=0$ for all pairs $(i,j)$.
Proof.
Step 1: Construct the Lyapunov Function. We construct a Lyapunov function representing the total disagreement in the system:
$V=\frac{1}{2}\sum_{i,j}\|u_{i}-u_{j}\|_{L^{2}}^{2}.$
This can be expressed compactly using the graph Laplacian $L$ of the coupling network as $V=\sum_{i}\int u_{i}^{T}Lu_{i}\,dx$.
Step 2: Analyze the Time Derivative. The time derivative $\frac{dV}{dt}$ can be shown to be negative semi-definite. The coupling term contributes a term proportional to $\lambda_{2}(L)V$, where $\lambda_{2}(L)$ is the algebraic connectivity (the second smallest eigenvalue) of the graph Laplacian $L$.
Step 3: Conclude Convergence. If the nonlinear terms $f_{i}$ are Lipschitz continuous, we can establish that $\frac{dV}{dt}\leq-\mu V$ for some $\mu>0$ that depends on $\lambda_{2}(L)$. By Grönwall’s inequality, $V(t)$ decays to zero exponentially, proving synchronization. ∎
Appendix B Appendix
This appendix provides supplementary material to support the main paper. Section A details the complete experimental setup, ensuring full reproducibility. Section B presents the unabridged experimental data and further analysis. Section C contains additional supplementary analyses, including computational costs and robustness checks.
B.1 Detailed Experimental Setup
B.1.1 Datasets and Tasks.
We evaluate our PDE-enhanced Transformer models on the Long Range Arena (LRA) benchmark Tay et al. (2020), which consists of five challenging tasks designed to assess the ability of sequence models to capture long-range dependencies:
  • ListOps: A synthetic task requiring hierarchical reasoning over sequences of up to 2,000 tokens, with 10 classes.
  • Text Classification: Document classification on IMDb movie reviews with sequences up to 4,000 tokens and 2 classes.
  • Text Retrieval: Matching queries with relevant documents, with sequences up to 4,000 tokens and 2 classes.
  • PathFinder: A visual reasoning task on $\times32$ images involving path connectivity detection, flattened to 1,024-token sequences, with 2 classes.
  • Image Classification: CIFAR-10 classification on $\times32$ images, flattened to 1,024-token sequences, with 10 classes.
B.1.2 Software and Hardware Environment.
All experiments were conducted on a high-performance computing cluster equipped with NVIDIA A100-80GB GPUs. The software stack included PyTorch 1.12.1, CUDA 11.3, and Python 3.9. Reproducibility was ensured by fixing the random seeds for Python (42), NumPy (0), and PyTorch (1) across all runs.
B.1.3 Model Hyperparameters.
All experiments use Transformer configurations optimized for A100-80GB GPUs. Configurations were tuned per task to maximize GPU utilization while maintaining training stability, as detailed in Table 8. A unified set of training hyperparameters, shown in Table 9, was used across all runs to ensure fair comparisons.
Table 8: Task-specific model configurations.
Task Dim Layers Heads Batch Size
ListOps 128 6 8 256
PathFinder 128 6 8 512
Text Cls. 128 6 8 128
Text Retrieval 128 4 8 64
Image Cls. 128 4 8 1024
Table 9: Unified training hyperparameters.
Hyperparameter Value
Optimizer AdamW ($\beta_{1}=0.9,\beta_{2}=0.98$)
Learning Rate $10^{-3}$ with linear warmup
(10,000 steps) & cosine decay
Training Epochs 50 with early stopping (patience=10)
Weight Decay $10^{-5}$
Gradient Clipping Max norm 1.0
Dropout 0.1
MLP Hidden Dim 512 (4x model dimension)
B.1.4 The detailed PDE operation
Each PDE operation implements the discrete diffusion equation:
$\frac{\partial u}{\partial t}=\alpha\nabla^{2}u$ (7)
where $\alpha$ is a learnable diffusion coefficient initialized to 0.1, and $\nabla^{2}$ is approximated using the finite difference method with reflective boundary conditions:
$\nabla^{2}u_{i}\approx u_{i+1}-2u_{i}+u_{i-1}$ (8)
The implementation uses padding with replicate mode to handle sequence boundaries, ensuring smooth transitions and avoiding edge artifacts. Layer normalization is applied after each PDE update to maintain training stability:
$u^{(t+1)}=\text{LayerNorm}(u^{(t)}+\alpha\nabla^{2}u^{(t)})$ (9)
B.2 Complete Experimental Data and Analysis
This section provides the unabridged data from our experiments, offering a more granular view of the results presented in the main paper and a deeper analysis of performance trends.
B.2.1 Main Experiment: PDE Integration Positions.
Figure 5: Seven PDE integration positions in Transformer architecture. (1) After Embedding, (2) After MLP, (3) Layer Diffusion, (4) Before LayerNorm, (5) In Attention, (6) Head Diffusion, and (7) After Attention. The performance analysis (right) shows that inserting the PDE diffusion layer after the embedding layer yields the largest improvement (+4.07 $p$ on LRA), while placing it after attention leads to performance degradation. Key insights highlight that early integration provides semantic regularization at the source and a stronger foundation for attention, whereas late integration can introduce destructive interference.
To complement the aggregated results in the main paper, we present the full performance data for our primary experiment comparing the seven PDE integration positions. Table 10 provides the complete performance breakdown for each model configuration on every LRA task. To better visualize these results, Figure 6 shows the per-task bar charts, while Figure 7 offers a heatmap for quick comparative analysis. These detailed results highlight task-specific sensitivities; for instance, the superiority of ‘PDE-After-Embed‘ is particularly pronounced on the PathFinder and Text Classification tasks, which heavily rely on spatial and semantic reasoning, respectively.
Table 10: Detailed performance comparison of different PDE integration positions on the LRA benchmark. Accuracy scores are reported for each task. Best performance per task is highlighted in bold.
Model ListOps Text Cls. Retrieval PathFinder Image Cls. Average
Baseline (Transformer) 0.3740 0.6480 0.8113 0.7017 0.3961 0.5862
After Embedding 0.3962 0.7029 0.8113 0.7943 0.4296 0.6269
After MLP 0.3896 0.6452 0.8233 0.7295 0.4053 0.5986
Layer Diffusion 0.3850 0.6600 0.8200 0.7100 0.4100 0.5970
Before LayerNorm 0.3896 0.6452 0.8113 0.7295 0.4053 0.5962
In Attention 0.3891 0.6842 0.8162 0.6872 0.3779 0.5909
Head Diffusion 0.3780 0.6550 0.8150 0.6942 0.4000 0.5884
After Attention 0.3740 0.6442 0.8112 0.5681 0.4109 0.5617
Figure 6: Detailed model performance across all five tasks in the Long Range Arena (LRA) benchmark. Each subplot shows the test accuracy for different PDE integration positions on a specific task.
Figure 7: Performance heatmap of all model configurations across the five LRA benchmark tasks. Warmer colors indicate higher performance.
B.2.2 Task-wise Improvement Analysis.
Table 11 further quantifies the differences observed in the previous section by detailing the absolute ($\Delta$) and relative (%) performance gains of each configuration over the baseline for every task. This granular analysis reveals task-specific sensitivities; for instance, ‘After Embedding‘ shows particularly strong gains on PathFinder (+13.20%), a task heavily reliant on spatial reasoning.
Table 11: Task-wise improvement analysis showing absolute ($\Delta$) and relative (%) gains over baseline for each PDE position.
Position ListOps Text Cls. Retrieval PathFinder Image Cls.
$\Delta$ % $\Delta$ % $\Delta$ % $\Delta$ % $\Delta$ %
After Embedding +0.0222 +5.94 +0.0549 +8.47 +0.0000 +0.00 +0.0926 +13.20 +0.0335 +8.46
After MLP +0.0156 +4.17 -0.0028 -0.43 +0.0120 +1.48 +0.0278 +3.96 +0.0092 +2.32
Layer Diffusion +0.0110 +2.94 +0.0120 +1.85 +0.0087 +1.07 +0.0083 +1.18 +0.0139 +3.51
Before LayerNorm +0.0156 +4.17 -0.0028 -0.43 +0.0000 +0.00 +0.0278 +3.96 +0.0092 +2.32
In Attention +0.0151 +4.04 +0.0362 +5.59 +0.0049 +0.60 -0.0145 -2.07 -0.0182 -4.59
Head Diffusion +0.0040 +1.07 +0.0070 +1.08 +0.0037 +0.46 -0.0075 -1.07 +0.0039 +0.98
After Attention +0.0000 +0.00 -0.0038 -0.59 -0.0001 -0.01 -0.1336 -19.04 +0.0148 +3.74
B.2.3 Ablation Study: Multi-Scale Dynamics.
This section provides the complete data for the multi-scale ablation study on the ListOps task. Table 12 presents the full numerical results, including the performance gain of the multi-scale approach over the best-performing single scale. The trends are visualized in Figure 8 (heatmap), Figure 9 (improvement plot), and Figure 10 (detailed bar charts), collectively demonstrating the consistent superiority of the adaptive multi-scale configuration.
Table 12: Complete multi-scale PDE ablation results on the ListOps task, including gain over the best single-scale performance.
Position Scale Configuration Best Multi-scale
Fast Medium Slow Multi-scale Single Gain
Baseline 0.3740
After Embedding 0.3960 0.3940 0.3990 0.4080 0.3990 +0.0090
After MLP 0.3900 0.3930 0.3910 0.4010 0.3930 +0.0080
Layer Diffusion 0.3850 0.3890 0.3910 0.3970 0.3910 +0.0060
Improvement over Baseline (0.3740):
After Embedding +0.0220 +0.0200 +0.0250 +0.0340 +0.0250 +0.0090
After MLP +0.0160 +0.0190 +0.0170 +0.0270 +0.0190 +0.0080
Layer Diffusion +0.0110 +0.0150 +0.0170 +0.0230 +0.0170 +0.0060
Figure 8: Heatmap of the multi-scale PDE ablation study on the ListOps task.
Figure 9: Improvement analysis of the multi-scale PDE ablation study on the ListOps task.
Figure 10: Detailed results of the multi-scale PDE ablation study on the ListOps task, broken down by PDE position.
B.2.4 Cross-Task Performance Consistency Analysis.
To assess the generalizability of our approach, we computed a cross-task performance correlation matrix (Table 13). Higher correlation values for our top models compared to the baseline indicate that the performance improvements are consistent across different task types.
Table 13: Cross-task performance correlation matrix for top-performing PDE positions. Higher values suggest more consistent performance improvements across different task types.
Task Pair After Embed. After MLP Layer Diff. Baseline $\Delta$ Correlation
ListOps $\leftrightarrow$ Text 0.82 0.65 0.71 0.58 +0.20
ListOps $\leftrightarrow$ Retrieval 0.45 0.73 0.68 0.41 +0.25
ListOps $\leftrightarrow$ PathFinder 0.89 0.72 0.76 0.67 +0.18
ListOps $\leftrightarrow$ Image 0.91 0.68 0.74 0.69 +0.16
Text $\leftrightarrow$ Retrieval 0.67 0.85 0.79 0.62 +0.19
Text $\leftrightarrow$ PathFinder 0.78 0.59 0.65 0.54 +0.16
Text $\leftrightarrow$ Image 0.83 0.61 0.67 0.58 +0.17
Retrieval $\leftrightarrow$ PathFinder 0.52 0.68 0.63 0.49 +0.14
Retrieval $\leftrightarrow$ Image 0.58 0.71 0.66 0.54 +0.13
PathFinder $\leftrightarrow$ Image 0.87 0.64 0.69 0.71 +0.12
B.3 Supplementary Analysis
B.3.1 Computational Cost Analysis.
The PDE Diffusion Layer is computationally efficient. For a sequence of length $N$ and dimension $D$, the additional cost is linear, $O(ND)$. In practice, this added approximately 5-15% to the total training time depending on the configuration, while delivering significant accuracy improvements. Memory overhead was negligible. Table 14 provides a detailed breakdown for key configurations on the ListOps task.
Table 14: Computational cost analysis on the ListOps task.
Config. Time OH (%) Mem. OH (%)
Baseline
After Embedding +5.6% +1.2%
After MLP +12.8% +2.4%
Multi-scale (Emb.) +18.4% +3.7%
B.3.2 Robustness and Sensitivity Analysis.
To ensure our results were not coincidental, we performed a limited evaluation with multiple random seeds on the top-performing positions for the ListOps task (Table 15). The results showed stable performance with small standard deviations, confirming the robustness of the observed improvements. We also tested sensitivity to the initialization of $\alpha$, finding that 0.1 provided a good balance between convergence speed and stability. Using ‘replicate‘ padding for boundary conditions was also found to be superior to zero padding, which could cause edge artifacts.
Table 15: Performance consistency across different random seeds on the ListOps task (mean $\pm$ std. dev.).
Position Accuracy
Baseline 0.3740 $\pm$ 0.0054
After Embedding 0.3962 $\pm$ 0.0042
After MLP 0.3896 $\pm$ 0.0039
Layer Diffusion 0.3850 $\pm$ 0.0035
B.3.3 Frequency Domain Analysis.
The success of multi-scale diffusion can be understood from a frequency domain perspective. The discrete diffusion operator acts as a low-pass filter. Different step sizes, $h$, correspond to different cutoff frequencies.
  • Fast scale ($h=1$) preserves high-frequency details.
  • Slow scale ($h=4$) emphasizes the low-frequency global structure.
The multi-scale approach creates a more balanced frequency response by combining these filters, allowing the model to capture a more comprehensive set of signal components. The transfer function for diffusion with step size $h$ is $H(\omega,h)=1-4\sin^{2}(\frac{\omega h}{2})$, and our multi-scale combination creates a more uniform response across the frequency spectrum, as shown in the energy distribution analysis in Table 16.
Table 16: Theoretical frequency band energy distribution.
Scale High Mid-H Mid-L Low
Fast ($h=1$) 0.81 0.40 1.75 2.81
Medium ($h=2$) 0.64 0.76 1.89 1.85
Slow ($h=4$) 0.42 1.25 1.67 0.93
Multi-scale 0.64 0.82 1.71 1.80
B.4 Core Mechanism and Theoretical Analysis
B.4.1 Theoretical Design Principles.
This seemingly simple formula embodies two key, principled design choices:
  1. 1.
    Update Form Based on Numerical Analysis. The residual form $X+\dots$ is not merely a conventional skip connection. In numerical analysis, it can be rigorously interpreted as a single-step forward Euler method for solving the diffusion equation $\frac{\partial u}{\partial t}=\alpha_{\text{coeff}}\nabla^{2}u$, where the time step $\Delta t$ is effectively absorbed into the learnable coefficient $\alpha$. According to Theorem 1.2 (Exponential Decay of the Gradient Norm), this update form guarantees that the system’s energy gradient converges at an exponential rate, providing a solid mathematical foundation for stability:
    $\left\|\frac{\delta E}{\delta u}(t+\Delta t)\right\|_{L^{2}}\leq\left\|\frac{\delta E}{\delta u}(t)\right\|_{L^{2}}e^{-\mu\alpha}$ (10)
    where $\mu$ is a positive definite constant of the system, ensuring the stability of the information propagation process.
  2. 2.
    Learnable Adaptive Diffusion Strength. The diffusion coefficient $\alpha$ is a learnable scalar parameter, which endows the model with a crucial adaptive capability. Based on Theorem 3.1 (Frequency Properties of Multi-Scale Filters), the learning process of $\alpha$ is, in fact, an optimization of the layer’s frequency response function:
    $\hat{\Delta}_{\text{adaptive}}(\omega)=\alpha\cdot\left(-\frac{4}{h^{2}}\sin^{2}\left(\frac{\omega h}{2}\right)\right)$ (11)
    The model can autonomously decide the required intensity of information smoothing at each layer based on the task and data, and can even effectively ”turn off” the layer by learning an $\alpha$ close to zero.
B.4.2 Theoretical Complementarity Analysis.
The PDE Diffusion Layer introduces a powerful inductive bias of local smoothness into the Transformer model. It forms a profound theoretical complementarity with the self-attention mechanism:
  • Self-Attention: Performs sparse, content-based global information retrieval.
  • PDE Diffusion: Performs dense, structure-based local information integration.
According to Theorem 2.1 (Polynomial Decay of the Heat Kernel), this complementarity is manifested in a fundamental difference in their information propagation patterns. For two positions at a distance of |xy||x-y|, their influence decay patterns are:
  • Attention Weights (Typical): $A(x,y)\sim e^{-c|x-y|}$ (Exponential decay)
  • Diffusion Kernel: $K_{t}(x,y)\sim|x-y|^{-1}$ (Polynomial decay)
Polynomial decay provides a stronger long-range connection capability, effectively supplementing the deficiencies of standard attention mechanisms in capturing ultra-long-range dependencies.
B.4.3 Inherent Positional Awareness.
The PDE Diffusion Layer provides the model with an inherent, structural sense of position that is orthogonal to traditional positional encodings. While standard Transformers rely on externally injected signals (e.g., sinusoidal encodings) to perceive order, our diffusion mechanism, through its intrinsic local stencil, tightly couples the representation of each token to its immediate neighbors, thus naturally encoding the topology and relative order of the sequence.

### v3

@@
-PDE-Transformer: A Continuous Dynamical Systems Approach to Sequence Modeling
+Where to Add PDE Diffusion in Transformers
 Abstract
-We propose PDE-Transformer, a novel sequence‐modeling paradigm that casts the forward pass of a Transformer as the numerical discretization of a continuous reaction–diffusion system derived from a variational energy functional.
-In our framework, token embeddings evolve under a partial differential equation whose nonlocal integral term models self-attention, local reaction term models feed-forward layers, diffusion term encodes positional smoothing, and a stability control term corresponds to layer normalization.
-From this unifying perspective, we design an Adaptive PDE Diffusion Layer—an efficient, learnable finite-difference stencil that enforces local smoothness in feature space with linear time complexity and complements self-attention’s global routing.
-Through a systematic theoretical analysis based on four pillars (stability, diffusion geometry, multi-scale dynamics, and component coupling), we derive principled guidelines for integrating the PDE layer at seven candidate points in the Transformer.
-Empirically, on the Long Range Arena benchmark, placing the layer immediately after embedding yields a 4.1 pp average accuracy gain over a strong baseline, and an adaptive multi-scale variant delivers further improvements.
-Our work thus offers a principled, lightweight mechanism to bolster long-range dependency modeling by harmonizing continuous PDE smoothing with discrete self-attention.
+Transformers achieve powerful content-based global routing via self-attention, but they do not explicitly encode a local geometric prior along the sequence axis, making the placement of locality-inducing modules largely empirical in prior hybrid designs.
+We study a simple, deterministic PDE diffusion layer—one explicit Euler step of 1D heat-equation smoothing implemented by a discrete Neumann Laplacian under a spectral stability budget—and ask a structural question: where should diffusion be inserted relative to attention?
+Our central claim is that diffusion and attention generally do not commute, so inserting the same local operator before versus after attention can induce qualitatively different behaviors.
+To formalize this, we develop a three-layer operator-theoretic framework that (i) provides unconditional guarantees for the diffusion subsystem (spectral non-expansiveness and monotone Dirichlet-energy dissipation under $\alpha<1/2$), (ii) derives compositional perturbation bounds linking insertion effects to representation roughness and downstream amplification, and (iii) uses diffusion–attention non-commutativity as a diagnostic for structural “double-mixing” conflicts.
+Guided by these predictions, we evaluate seven canonical insertion positions on Long Range Arena and find that early diffusion acts as effective pre-regularization, improving average accuracy by +4.1+4.1 points when applied after embedding, whereas post-attention diffusion degrades performance by 2.5-2.5 points, consistent with the predicted conflict.
+We further show that a multi-scale diffusion variant yields consistent gains under the same global stability budget.
+Beyond this specific module, our analysis provides a transferable template for reasoning about local–global compositions in sequence models by cleanly separating provable guarantees, compositional bounds, and mechanistic diagnostics.
+Transformers, locality, PDE diffusion, operator theory, spectral stability
 1 Introduction
-Since its inception, the Transformer architecture has achieved revolutionary breakthroughs across diverse fields, from natural language processing and computer vision to protein folding, owing to its powerful self-attention mechanism.
-However, as applications scale from short texts to ultra-long sequences—such as document-level dialogues or whole-genome sequences—the standard Transformer reveals two fundamental bottlenecks.
-First, its computational and memory complexity, which scales quadratically with sequence length $L$ ($O(L^{2})$), becomes prohibitively expensive.
-Second, its purely content-driven global interaction mechanism lacks explicit modeling of local geometric structure, hindering its ability to effectively capture long-range dependencies.
-Although variants like sparse attention and low-rank approximations offer patchwork optimizations, they remain engineering modifications to a discrete computational graph, failing to transcend the limitations of the discrete paradigm or provide a unified theoretical framework for their design.
-To overcome this impasse, we propose a return to the first principles of physics, reframing the process of sequence modeling as a continuous variational dynamical system.
-In this framework, token embeddings do not undergo discrete layered transformations but rather evolve within an energy field governed by three fundamental forces: local diffusion, nonlinear reaction, and nonlocal coupling.
-We establish a one-to-one correspondence between these forces and the core components of the Transformer: the nonlinear reaction maps to the feed-forward network (FFN), and the nonlocal coupling maps to the self-attention mechanism.
-This perspective strikingly reveals a structural deficiency in the standard Transformer: the absence of the crucial diffusion term.
-In physical systems, this term is responsible for penalizing sharp variations and imposing local smoothness, which is key to forming stable, ordered structures.
-Building on this core finding, we design a lightweight, plug-and-play Adaptive PDE Diffusion Layer with linear complexity.
-This module discretizes a reaction-diffusion equation via a learnable finite-difference method, introducing a structured local smoothness inductive bias into the model.
-It forms a natural complementarity with self-attention: the PDE diffusion layer is responsible for reinforcing local geometric consistency, while self-attention focuses on capturing global content-based associations.
-To validate this paradigm, we systematically investigate the impact of integrating the PDE diffusion layer at seven distinct points within the Transformer architecture.
-Experiments on the challenging Long Range Arena (LRA) benchmark provide compelling evidence for our theory: placing the PDE layer immediately after the initial token embeddings and before the first Transformer block yields the most significant performance gains.
-This configuration achieves an average accuracy improvement of 4.1 percentage points over a strong baseline, with a multi-scale variant delivering further improvements.
-This finding uncovers a core mechanism: before global, sparse attention interactions can be effective, the raw semantic space must first undergo local, dense, structured smoothing.
-The main contributions of this work are fourfold.
-First, we introduce a new theoretical paradigm that unifies sequence modeling as a continuous reaction–diffusion dynamical system derived from variational principles, offering a novel physical lens to understand and improve Transformer architectures.
-Second, we design the Adaptive PDE Diffusion Layer as a plug-and-play, linear-complexity module that significantly enhances the model’s ability to capture local structure at negligible computational cost.
-Third, we conduct a systematic empirical study to identify the optimal integration point for such a local smoothing mechanism within the Transformer.
-Finally, we reveal the profound complementarity between explicit local geometric smoothing and global content-based aggregation, providing principled guidance for building more robust and efficient long-sequence models.
-The remainder of this paper is organized as follows.
-Section 2 reviews prior work on efficient Transformers and continuous-time models.
-Section 3 details our theoretical framework, deriving the PDE-Transformer from first principles.
-Section 4 presents our comprehensive experimental validation on the LRA benchmark.
-Finally, Section 5 concludes the paper and discusses future directions.
+1.1 Motivation
+Transformers have become the dominant architecture for sequence modeling, largely due to self-attention’s ability to perform content-based global routing (Vaswani et al., 2017).
+By enabling interactions between arbitrary token pairs, attention efficiently captures long-range dependencies and global semantic structure.
+However, standard Transformers lack an explicit local geometric prior along the sequence dimension: they do not directly encourage neighboring token representations to vary smoothly or preserve local coherence.
+While attention can learn locality implicitly, it is not encoded architecturally and may be brittle under distribution shifts (e.g., different sequence lengths).
+This gap has motivated hybrid designs that inject locality through convolutional layers (Wu et al., 2019; Dai et al., 2020), sparse patterns (Beltagy et al., 2020), or other structured inductive biases.
+Yet a surprisingly basic question remains under-theorized: where should a local operator be placed relative to attention?
+In practice, placement is often treated as a hyperparameter and tuned by validation performance.
+In this work, we ask whether placement can be predicted from structure rather than guessed from sweeps: can operator theory explain when local smoothing helps and when it hurts?
+Our central observation is simple but consequential: local smoothing and content-based global routing are, in general, non-commuting operations.
+Let $\mathcal{S}$ denote a local PDE diffusion (smoothing) operator and $\mathcal{A}$ an attention-based routing operator.
+Typically, $\mathcal{S}\circ\mathcal{A}\neq\mathcal{A}\circ\mathcal{S}$, and the commutator $\mathcal{S},\mathcal{A}]=\mathcal{S}\mathcal{A}-\mathcal{A}\mathcal{S}$ captures this incompatibility.
+This has direct architectural implications: applying PDE diffusion before attention can pre-regularize a rough representation space and provide a smoother manifold for routing; applying PDE diffusion after attention can instead blur structures just created by non-local mixing, inducing a harmful double-mixing effect.
+1.2 Our Approach: A Three-Layer Operator-Theoretic Framework
+To formalize these intuitions, we develop a systematic operator-theoretic framework for analyzing interactions between locality and attention in Transformers.
+The framework is organized into three conceptual layers, each addressing a different aspect of the problem.
+Layer I: Intrinsic stability (provable guarantees).
+We first study the PDE diffusion operator in isolation.
+Using a discrete Laplacian with a strict spectral budget constraint ($\alpha<1/2$), we establish unconditional guarantees for the PDE diffusion subsystem alone.
+Specifically, we show that the operator is spectrally non-expansive, with strict contraction on non-constant modes, and that it induces monotone dissipation of the discrete Dirichlet energy.
+These results formalize PDE diffusion as a stable geometric low-pass filter that reduces local roughness independently of the Transformer’s non-convex components, thereby providing a mathematically grounded stability anchor.
+Takeaway: under a simple spectral budget, PDE diffusion is a certified non-expansive local smoother.
+Layer II: Compositional analysis (perturbation bounds).
+We then analyze how this stable PDE diffusion operator interacts with the rest of the Transformer when composed at different locations.
+Treating downstream modules—such as LayerNorm, attention, and MLPs—as locally Lipschitz mappings, we derive perturbation propagation bounds showing that the effect of inserting PDE diffusion at a given position depends on two factors: the geometric roughness of the representation at the insertion point, and the sensitivity of subsequent layers to perturbations.
+This analysis explains why the same PDE diffusion operator can be beneficial when applied early, where representations are rough and regularization is helpful, yet detrimental when applied late, where representations are already smooth and additional smoothing becomes unnecessary or harmful.
+Takeaway: PDE diffusion helps most where representations are rough and downstream amplification is limited—typically early in the network.
+Layer III: Non-commutativity diagnostics (design guidelines).
+Finally, we turn to the interaction between PDE diffusion and attention.
+We formalize their incompatibility through the commutator $\mathcal{S},\mathcal{A}](X)$, which measures the representational discrepancy between pre-attention and post-attention diffusion.
+Under a local linearization of attention, this conflict scales with the mismatch between the local geometric structure enforced by PDE diffusion and the non-local routing patterns produced by attention.
+This perspective yields concrete, testable design guidelines: PDE diffusion should be applied early as a form of semantic manifold pre-regularization; post-attention PDE diffusion should be avoided due to structural conflict; and multi-scale PDE diffusion, when constrained by a global spectral budget, can improve robustness by covering multiple frequency bands without sacrificing stability.
+Takeaway: because PDE diffusion and attention generally do not commute, post-attention smoothing can erase structures created by global routing.
+1.3 Contributions
+This paper makes four main contributions.
+First, we develop a three-layer operator-theoretic framework that separates provable intrinsic guarantees from compositional perturbation analysis and architectural design heuristics, clarifying what can and cannot be established rigorously.
+Second, we derive principled architectural guidelines for inserting locality-inducing modules into Transformers, predicting that early-stage insertion should systematically outperform post-attention insertion due to fundamental operator incompatibility.
+Third, we provide controlled empirical validation on the Long Range Arena benchmark (Tay et al., 2020b), systematically evaluating seven insertion positions.
+The results align closely with theory: PDE diffusion after embedding yields a +4.1+4.1 percentage point improvement, while PDE diffusion after attention leads to a 2.5-2.5 point degradation.
+Finally, beyond the specific PDE diffusion operator studied here, we demonstrate how operator-theoretic reasoning can guide the design of hybrid architectures, offering a transferable methodology applicable to other local–global composition problems, such as convolution–attention or state-space–attention hybrids.
+1.4 Roadmap
+The remainder of the paper is organized as follows.
+Section 2 reviews related hybrid architectures and theoretical perspectives.
+Section 3 introduces the operator-theoretic framework and defines the PDE diffusion layer precisely.
+Section 4 presents the three-layer theoretical analysis, including intrinsic stability guarantees and compositional perturbation bounds.
+Section 5 empirically validates the resulting architectural guidelines on Long Range Arena.
+Section 6 concludes.
 2 Related Work
-We situate the PDE-Transformer within three major research streams: (i) efficiency enhancements to discrete Transformers, (ii) continuous-time sequence modeling, and (iii) neural networks as PDE solvers.
-Together, these perspectives highlight both the progress and the remaining gaps that motivate our work.
-2.1 Discrete Efficiency Enhancements
-The quadratic complexity of the vanilla Transformer Vaswani et al.
-(2017) has motivated extensive efforts to improve scalability.
-One class of methods reduces attention cost by introducing fixed or learned sparsity, such as local windows combined with global tokens Beltagy et al.
-(2020); Zaheer et al.
-(2020) or strided/dilated patterns Child et al.
-(2019).
-While effective, these approaches impose pre-defined structures that may not align with data and risk bottlenecks in capturing long-range dependencies.
-A second class leverages low-rank or kernel approximations: Linformer Wang et al.
-(2020) projects Key and Value matrices into lower-dimensional spaces, while Performer Choromanski et al.
-(2021) approximates softmax attention with kernel features.
-These methods reduce complexity but rely on restrictive low-rank assumptions that can discard fine-grained information in the long-tail spectrum.
-A third family introduces recurrence and memory, such as Transformer-XL Dai et al.
-(2019), which caches hidden states across segments to extend context length.
-This alleviates quadratic cost but introduces challenges in compressing historical states and handling discontinuities at segment boundaries.
-Overall, these methods optimize the discrete computation graph—deciding “who attends to whom”—without altering the fundamental discrete paradigm (see Table 1).
-Table 1: Taxonomy of efficient Transformer architectures. These methods improve scalability within a discrete framework but do not change the underlying paradigm.
-Approach Key Examples Core Idea Limitation
-Sparsity BigBird Zaheer et al. (2020), Longformer Beltagy et al. (2020) Restrict attention to sparse patterns Rigid information flow; may miss dependencies
-Low-Rank / Kernel Linformer Wang et al. (2020), Performer Choromanski et al. (2021) Approximate the full attention matrix Low-rank assumption; loss of high-frequency info
-Recurrence / Memory Transformer-XL Dai et al. (2019) Segment sequence and reuse past states State compression bottleneck; boundary effects
-2.2 Continuous-Time Models
-A more fundamental line of research interprets neural networks as discretizations of continuous-time systems.
-Neural ODEs Chen et al.
-(2018) recast residual networks as solutions of ODEs, treating depth as continuous and enabling benefits such as memory-efficient training Gholami et al.
-(2019) and irregular time-series modeling Li et al.
-(2020).
-However, these methods continuous-ize the depth dimension rather than the sequence.
-State-Space Models (SSMs) directly target the sequence dimension. They model dynamics with a linear ODE:
-$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t)+\mathbf{D}x(t),$ (1)
-where $x(t)$ is the input, $h(t)$ the latent state, and $y(t)$ the output.
-The S4 model Gu et al.
-(2022) demonstrated that by structuring $\mathbf{A}$, one can discretize this system into an efficient convolutional form.
-Recent advances like Mamba Gu & Dao (2023) incorporate selection mechanisms and input-dependent parameters, surpassing linear time-invariant constraints and achieving strong performance.
-SSMs thus represent the most competitive continuous-time alternative to Transformers.
-However, they are governed by ODEs, which describe the temporal evolution of state vectors.
-Our PDE formulation instead models a field $u(x,t)$, introducing explicit spatial derivative terms (e.g., $\nabla^{2}u$) and enabling spatiotemporal coupling along the sequence dimension (Table 2).
-Table 2: Comparison of continuous-time modeling paradigms. Unlike ODE-based models, our PDE approach introduces explicit spatiotemporal coupling.
-Paradigm Continuous Dim. Equation Mechanism Limitation
-Neural ODE Chen et al. (2018) Depth ODE Continuous layers Models depth, not sequence
-SSM (Mamba) Gu & Dao (2023) Sequence ODE Temporal evolution + selection Lacks explicit spatial coupling
-PDE-Transformer (Ours) Sequence PDE Spatiotemporal coupling Models direct interaction via $\nabla^{2}u$
-2.3 Neural Networks as PDE Solvers
-A parallel literature connects neural networks and PDEs but with inverse objectives.
-Physics-Informed Neural Networks (PINNs) Raissi et al.
-(2019) parameterize PDE solutions with neural networks, enforcing boundary conditions and PDE residuals during training.
-Neural Operators such as the Fourier Neural Operator (FNO) Li et al.
-(2021) extend this idea by learning resolution-invariant mappings from input functions to PDE solutions, effectively serving as universal solvers.
-In contrast, our work does not seek to solve externally-given PDEs.
-Instead, we adopt PDE principles—diffusion, waves, and reaction—as internal design guidelines for the information dynamics of sequence models.
-2.4 Synthesis
-In summary, discrete efficiency methods enhance scalability without altering the static computation graph; continuous-time models, particularly SSMs, provide strong alternatives but are ODE-based and lack explicit spatial coupling; and neural PDE solvers apply PDEs as external constraints rather than intrinsic dynamics.
-This landscape reveals an open frontier: applying PDEs as first-principles foundations for neural sequence architectures.
-Our PDE-Transformer advances this direction by formalizing spatiotemporal coupling as a core design principle for long-range modeling.
-3 Theory and Method: Unified Variational Dynamical System and Adaptive Diffusion
-To address the dual bottlenecks of quadratic complexity and the lack of local geometric modeling in Transformers, we reframe sequence modeling as a variational dynamical system from first principles.
-In this view, representations evolve under local diffusion, nonlinear reaction, and nonlocal coupling forces.
-This reveals a structural deficiency in standard Transformers—the missing diffusion term—which we remedy with a practical and efficient adaptive PDE diffusion layer.
-In this section, we derive the governing dynamics from an energy functional, map its components to Transformer modules, and instantiate the theory as a trainable layer, further proposing a principled integration strategy and discussing its multiscale spectral properties.
-3.1 Unified Variational Principle
-(a) Unified variational framework: energy functional decomposition and gradient-flow PDE.
-(b) Architecture mapping: PDE components vs. standard Transformer, showing the missing diffusion term and our proposed PDE diffusion layer.
-Figure 1: Theory framework of PDE-Transformer.
-Two stacked panels show (a) the unified variational formulation and corresponding PDE, and (b) the architectural mapping that highlights the missing diffusion component in standard Transformers.
-We denote the continuous representation of the sequence as $u(x,t)$, where $x\in\Omega$ represents the position and $t$ the “evolution time” corresponding to network depth.
-The global energy functional is defined as
-$E[u]$ $=\int_{\Omega}\Big(\tfrac{\alpha}{2}|\nabla u|^{2}+F(u)\Big)dx+\frac{\beta}{2}\iint_{\Omega\times\Omega}K(x,y)[u(x)-u(y)]^{2}dxdy.$ (2)
-It consists of three parts: a local tension term $|\nabla u|^{2}$ that penalizes sharp variations to maintain local consistency; a nonlinear potential $F(u)$ encoding task objectives; and a long-range coupling $K(x,y)$ establishing content-dependent global interactions.
-Theorem 3.1 (Unified Dynamical Equation).
-The gradient flow of $E[u]$ is
-$\frac{\partial u}{\partial t}=\alpha\Delta u-F^{\prime}(u)-\beta\mathcal{L}_{K}[u],\qquad\mathcal{L}_{K}[u](x)=\int K(x,y)[u(x)-u(y)]dy.$ (3)
-Each term corresponds directly to a Transformer module. The standard Transformer realizes reaction and nonlocal coupling but ignores diffusion, limiting explicit geometric modeling and stability.
-Table 3: Physical and functional correspondence of PDE-Transformer components.
-Mathematical Term Physical Process Module Function Information Pattern
-$\alpha\Delta u$ Diffusion PDE Diffusion Layer (Ours) Regularizes locally Geometric, dense
-$F^{\prime}(u)$ Reaction Feedforward Network Pointwise transformation Local, independent
-$\beta\mathcal{L}_{K}[u]$ Nonlocal coupling Self-Attention Aggregates globally Sparse, content-driven
-3.2 Diffusion Dynamics and Stability
-To align theory with implementation, we adopt Neumann (reflective) boundary conditions:
-$(\mathbf{\Delta}_{N}X)_{i}=\begin{cases}X_{2}-X_{1},&i=1,\\ X_{i-1}-2X_{i}+X_{i+1},&2\leq i\leq L-1,\\ X_{L-1}-X_{L},&i=L.\end{cases}$ (4)
-This preserves boundary information and is compatible with discrete cosine transform (DCT) analysis.
-Theorem 3.2 (Spectrum).
-The eigenvalues of $\mathbf{\Delta}_{N}$ are $\lambda_{k}=-4\sin^{2}(\pi k/(2L))$ with cosine eigenvectors.
-Corollary 3.3 (CFL Condition).
-For the explicit iteration $X^{(k+1)}=X^{(k)}+\alpha\mathbf{\Delta}_{N}X^{(k)}$, stability requires $\alpha<\tfrac{1}{2}$.
-Theorem 3.4 (Lyapunov Monotonicity).
-Under the CFL condition, the Dirichlet energy $V(X)=X^{\top}(-\mathbf{\Delta}_{N})X$ decreases monotonically, ensuring stability and improving the optimization landscape.
-3.3 Information-Theoretic Integration
-Consider a forward chain $X_{0}\to X_{1}\to\cdots\to X_{N}\to Y$. For any smoothing operator $\mathcal{S}$, define the information retention:
-$\rho(d)=\mathbb{E}\left[\frac{I(\mathcal{S}(X^{(d)});Y)}{I(X^{(d)};Y)}\right],\qquad\rho(1)\geq\rho(2)\geq\cdots\geq\rho(N).$ (5)
-This shows that earlier diffusion preserves more task-relevant information.
-We further compare seven candidate insertion points using the value function $V_{i}=w_{I}I_{i}-w_{D}D_{i}-w_{C}C_{i}$, where $I_{i}$ denotes information retention, $D_{i}$ distortion, and $C_{i}$ computational cost.
-Table 4: Comparison of candidate insertion points for diffusion layers.
-Pos. Mechanism Property Info Dist. Cost Rank
-1 After Embedding $\|X_{0}^{(s)}-X_{0}\|=O(\alpha)$ High Low Low 1
-2 After MLP Local proximal step High Low Low 2
-3 Between Layers Gradient flow improvement Mid-High Mid Mid 3
-4 Before LN Non-commutativity Mid Mid Low 4
-5 Inside Attention Breaks normalization Mid-Low Mid-High Mid 5
-6 Between Heads Loss of diversity Low High Mid 6
-7 After Attention Breaks sparsity Low High Low 7
-Overall Ranking: After embedding >> After MLP >> Between layers >> Before LN >> Inside attention >> Between heads >> After attention.
-3.4 Multi-Scale Spectral Complementarity
-Figure 2: Frequency domain analysis of the multi-scale diffusion mechanism.
-(Left) The transfer function $H(\omega)$ for different diffusion scales.
-Single-scale diffusions (Fast, Medium, Slow) act as low-pass filters with different cutoff frequencies.
-(Right) The energy distribution across four frequency bands.
-The multi-scale approach (green, dashed line) achieves a more balanced energy distribution across the entire frequency spectrum compared to any single-scale method, enabling it to capture a richer set of signal components from both global trends (low frequency) and local details (high frequency).
-For diffusion with step size $h$, the frequency response is
-$H_{h}(\omega)=1-4\alpha\sin^{2}\!\Big(\tfrac{\omega h}{2}\Big).$ (6)
-Small steps ($h=1$) preserve high-frequency details, while larger steps ($h=2,4$) emphasize mid-to-low frequencies.
-Mixing scales covers the entire spectrum; a practical initialization uses weights $1:0.6:0.3$.
-Diffusion is geometry-driven and dense, with effective radius proportional to $\sqrt{t}$, while attention is content-driven and sparse, connecting arbitrary positions.
-Their combination captures both local geometry and global dependency.
-3.5 Adaptive PDE Diffusion Layer
-The proposed module incorporates multi-scale diffusion ($[1,2,4]$), learnable coefficients per scale and channel, and CFL-enforced constraints $\alpha<0.5$ at runtime.
-Its complexity is $O(Ld)$ for a single scale or $O(KLd)$ for multiple scales, significantly lower than attention $O(L^{2}d)$.
-Theory and information analysis indicate that the best insertion point is after the embedding layer, before the first Transformer block, maximizing information retention and stability while improving optimization.
-Summary.
-We unify sequence modeling into a reaction–diffusion–nonlocal coupling system, derive its integro-PDE gradient flow, and identify the missing diffusion term in Transformers.
-Spectral and stability analysis motivates an efficient adaptive PDE diffusion layer, with its optimal position after embedding.
-This principled module augments global coupling with explicit local modeling, laying the foundation for subsequent experiments.
-applications.
+Our work sits at the intersection of (i) long-sequence Transformers, (ii) locality-inducing inductive biases, and (iii) stability- and operator-motivated architectural analysis.
+We position our contribution along three axes: we do not change attention’s routing rule or complexity (unlike efficient-attention variants), we use PDE diffusion as a fixed-form, spectrally-budgeted local prior (not to be confused with generative diffusion models), and we provide a structural explanation for insertion-position sensitivity (often treated empirically in hybrid designs).
+2.1 Long-Sequence Transformers
+The quadratic complexity of self-attention has motivated extensive work on efficient long-sequence Transformers.
+Representative directions include sparse patterns (e.g., Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), Sparse Transformers (Child et al., 2019), Reformer (Kitaev et al., 2020)), low-rank or kernelized approximations (e.g., Performer (Choromanski et al., 2020), Linear Transformers (Katharopoulos et al., 2020), Linformer (Wang et al., 2020), Nyströmformer (Xiong et al., 2021)), and memory-optimized exact attention (e.g., FlashAttention (Dao et al., 2022)).
+LRA (Tay et al., 2020b, a) further highlights that efficiency improvements do not automatically translate into robust long-context generalization.
+Our work is orthogonal: we keep attention intact and instead study how to inject an explicit local geometric prior via a lightweight PDE diffusion operator.
+Because the PDE diffusion layer is linear-time, it can be combined with sparse/linear attention mechanisms as a complementary design component.
+2.2 Diffusion, PDEs, and Neural Operators
+PDE- and diffusion-inspired views have a long history in deep learning, including stability-aware architectures via dynamical-systems discretizations (Haber & Ruthotto, 2017; Ruthotto & Haber, 2018) and classical diffusion for denoising/smoothing (Weickert, 1998).
+More recently, neural operator learning aims to approximate families of PDE solution operators (Li et al., 2020), and generative diffusion models treat diffusion as a stochastic generative process (Ho et al., 2020).
+Our use of PDE diffusion differs in both goal and modeling choice.
+We use a fixed-form discrete Laplacian with a strict spectral budget to inject a local smoothness prior with unconditional stability guarantees, rather than learning an operator approximator or modeling a generative diffusion process.
+The resulting module is a deterministic architectural regularizer that targets the sequence dimension and integrates naturally with Transformers.
+2.3 Spectral Stability and Regularization in Deep Networks
+Stability and spectral control have been recognized as central issues in deep learning, particularly for very deep or long-context models.
+Spectral normalization (Miyato et al., 2018) and Lipschitz-constrained architectures aim to control gradient explosion and improve robustness.
+Layer normalization (Ba et al., 2016) and residual connections (He et al., 2015) provide fundamental stability mechanisms in deep networks.
+In Transformers, prior work has analyzed the instability of attention maps and the amplification of high-frequency components across layers.
+Our work contributes to this line by introducing a simple yet effective mechanism for spectral control along the sequence dimension.
+By leveraging the discrete Laplacian with a strict spectral budget, we guarantee non-expansiveness and monotone Dirichlet energy dissipation for the diffusion subsystem.
+Importantly, our analysis does not rely on global Lipschitz assumptions for the full Transformer; instead, it isolates what can be proven unconditionally and studies interactions with attention through controlled perturbation and non-commutativity.
+2.4 Position Sensitivity and Hybrid Architectures
+Hybrid architectures that combine local and global operations are increasingly common, including convolution–attention hybrids (Wu et al., 2019; Dai et al., 2020) and alternatives to attention such as structured state space models (Gu et al., 2021; Gu & Dao, 2023).
+In many such designs, the placement of local operators relative to global mixing is decided empirically.
+Our contribution is a principled account of this placement sensitivity.
+By framing diffusion and attention as generally non-commuting operators, we explain why pre-routing diffusion can act as beneficial manifold regularization while post-routing diffusion can introduce a destructive double-mixing effect.
+This operator-theoretic lens is not specific to diffusion and can serve as a transferable template for reasoning about other local–global compositions.
+3 Operator Framework and Diffusion Layer Design
+3.1 Scope of Theoretical Guarantees
+This section introduces an operator-theoretic framework centered on a PDE diffusion layer and clarifies the precise scope of its theoretical guarantees.
+Our analysis is intentionally structured into three conceptual layers, each corresponding to a different level of rigor and assumption.
+At the first level, we establish intrinsic stability properties of the diffusion operator itself.
+These include spectral non-expansiveness and monotone energy dissipation, which hold unconditionally and independently of all other Transformer components.
+These results are provable without assumptions on attention, normalization, or nonlinear activations.
+At the second level, we study how diffusion interacts compositionally with backbone modules such as LayerNorm, self-attention, and MLP blocks.
+This analysis relies on local Lipschitz assumptions and yields perturbation propagation bounds that explain why the effect of diffusion depends strongly on its insertion position within the network.
+At the third level, we provide a diagnostic and heuristic analysis based on the non-commutativity between diffusion and attention operators.
+This layer depends on a local linearization of attention and does not claim strict guarantees.
+Instead, it offers principled architectural guidance by quantifying geometric conflict between local smoothing and global routing.
+This layered structure explicitly separates what is rigorously provable from what requires additional assumptions, thereby avoiding overstatement of theoretical claims.
+3.2 Core Operator Definitions
+We consider a sequence of length $L$ with hidden dimension $d$. Let $X\in\mathbb{R}^{L\times d}$ denote a layer representation, where $X_{i}\in\mathbb{R}^{d}$ is the feature vector at position $i$.
+Boundary conditions.
+To preserve information at sequence boundaries, we impose Neumann (zero-flux) boundary conditions via endpoint replication,
+$x_{0}=x_{1},\qquad x_{L+1}=x_{L}.$
+Discrete Neumann Laplacian.
+For a vector $x\in\mathbb{R}^{L}$, the discrete Neumann Laplacian $\Delta_{N}$ acts elementwise as
+$(\Delta_{N}x)_{i}=\begin{cases}x_{2}-x_{1},&i=1,\\ x_{i-1}-2x_{i}+x_{i+1},&2\leq i\leq L-1,\\ x_{L-1}-x_{L},&i=L.\end{cases}$ (1)
+For matrix inputs, the operator is applied channel-wise along the sequence dimension, i.e., $(\Delta_{N}X)_{\cdot,j}=\Delta_{N}(X_{\cdot,j})$.
+Single-scale diffusion.
+The minimal diffusion module is defined as
+$\mathcal{S}_{\alpha}(X)=(I+\alpha\Delta_{N})X,\qquad 0\leq\alpha<\tfrac{1}{2}.$
+The constraint $\alpha<\tfrac{1}{2}$ acts as a stability budget, analogous to the CFL condition in explicit diffusion schemes, and guarantees spectral non-expansiveness and energy dissipation.
+Multi-scale diffusion.
+To accommodate heterogeneous feature lengths, we extend the operator to a multi-scale formulation by combining $K$ strided Laplacians with step sizes $h_{k}$:
+$\mathcal{S}_{\boldsymbol{\alpha}}(X)=X+\sum_{k=1}^{K}\alpha_{k}\Delta_{N,h_{k}}X.$
+A global budget constraint,
+$\sum_{k=1}^{K}\alpha_{k}<\tfrac{1}{2},$
+ensures that the same stability guarantees are preserved.
+3.3 Systematic Insertion Design Space
+We define a systematic design space consisting of seven canonical insertion strategies, ranging from early-stage pre-regularization to late-stage post-routing smoothing.
+These positions are summarized in Table 1.
+Table 1: Canonical insertion strategies for PDE diffusion layers, spanning the full two-column width.
+ID Strategy Composition Theoretical Expectation
+P1 After-Embedding $X\leftarrow\mathcal{S}(\mathrm{Embed}(\cdot))$ Optimal: semantic manifold pre-regularization
+P2 After-MLP $X\leftarrow\mathcal{S}(X+\mathrm{MLP}(X))$ Moderate: partial regularization
+P3 Layer-Diffusion $X_{\ell+1}\leftarrow\mathcal{S}(\mathrm{Block}(X_{\ell}))$ Moderate: inter-layer smoothing
+P4 Before-LN $X\leftarrow\mathrm{LN}(\mathcal{S}(X))$ Moderate: interaction with normalization
+P5 In-Attention $\mathrm{Attn}(Q,K,V)\!\rightarrow\!\mathrm{Attn}(Q,K,\mathcal{S}(V))$ Unclear: internal intervention
+P6 Head-Diffusion $H\leftarrow\mathcal{S}_{\text{head}}(H)$ Unclear: cross-head exchange
+P7 After-Attention $X\leftarrow\mathcal{S}(X+\mathrm{Attn}(X))$ Worst: double-mixing conflict
+The central contrast lies between early and late insertion.
+Strategy P1 smooths representations before global routing, allowing attention to aggregate information on a less noisy semantic manifold.
+In contrast, P7 applies local smoothing after non-local mixing, potentially blurring structures that attention has just constructed.
+3.4 Implementation Details
+The PDE diffusion operator has linear complexity in both time and memory. A single-scale diffusion requires $\mathcal{O}(Ld)$ operations, while the multi-scale variant scales as $\mathcal{O}(KLd)$.
+To enforce the global stability budget exactly during training without manual clipping, we reparameterize the diffusion weights as
+$\alpha_{k}=\tfrac{1}{2}\,\sigma(\eta)\,\mathrm{softmax}(\boldsymbol{\theta})_{k},$
+where $\eta\in\mathbb{R}$ and $\boldsymbol{\theta}\in\mathbb{R}^{K}$ are unconstrained learnable parameters.
+Depending on architectural choice, diffusion can be applied either directly or followed by normalization,
+$\text{Out}=\mathcal{S}(X),\qquad\text{or}\qquad\text{Out}=\mathrm{LN}(\mathcal{S}(X)).$
+3.5 Analysis Quantities: Roughness and Conflict Diagnostics
+Local roughness.
+We quantify local geometric irregularity along the sequence dimension using the discrete Dirichlet energy,
+$\mathcal{E}(X)=\tfrac{1}{2}\langle X,-\Delta_{N}X\rangle=\tfrac{1}{2}\sum_{j=1}^{d}\sum_{i=1}^{L-1}\bigl(X_{i+1,j}-X_{i,j}\bigr)^{2}.$
+This quantity measures how rapidly representations vary between neighboring tokens.
+Geometric conflict.
+To diagnose incompatibility between diffusion and attention, we define the diffusion–attention commutator,
+$\mathcal{S},\mathcal{A}](X)=\mathcal{S}(\mathcal{A}(X))-\mathcal{A}(\mathcal{S}(X)).$
+This operator measures the representational discrepancy between pre-attention and post-attention diffusion.
+In this work, we use it as a conceptual diagnostic for architectural compatibility (Layer III), rather than as a measured training-time signal.
+4 Operator-Theoretic Analysis
+This section develops an operator-theoretic account of how locality-inducing PDE diffusion interacts with self-attention.
+To avoid overclaiming, we separate results into three layers: (i) unconditional properties of the PDE diffusion operator in isolation; (ii) compositional perturbation bounds under standard local regularity assumptions for Transformer modules; and (iii) a diagnostic interpretation of diffusion–attention non-commutativity based on a local linearization of attention.
+Full proofs are deferred to Appendix B.
+4.1 Layer I: Intrinsic Stability of Diffusion
+We begin with properties that hold independently of the Transformer backbone.
+Let $\Delta_{N}\in\mathbb{R}^{L\times L}$ denote the discrete Neumann Laplacian (Eq.
+(1)), and define the single-step diffusion operator
+$\mathcal{S}_{\alpha}\triangleq I+\alpha\Delta_{N},\qquad 0\leq\alpha<\tfrac{1}{2},$ (2)
+applied channel-wise to $X\in\mathbb{R}^{L\times d}$.
+Lemma 4.1 (Spectrum of the Discrete Neumann Laplacian).
+The matrix $\Delta_{N}$ is real symmetric and negative semi-definite. Its eigenvalues are
+$\lambda_{k}=-4\sin^{2}\!\left(\frac{\pi k}{2L}\right),\quad k=0,1,\ldots,L-1,$ (3)
+where $\lambda_{0}=0$ (constant mode) and $\lambda_{k}\in[-4,0]$ for all $k$.
+Theorem 4.2 (Spectral Non-Expansiveness and Strict Contraction).
+For any $X\in\mathbb{R}^{L\times d}$ and $0\leq\alpha<\tfrac{1}{2}$,
+$\lVert\mathcal{S}_{\alpha}\rVert_{2}\leq 1,\qquad\lVert\mathcal{S}_{\alpha}(X)\rVert_{F}\leq\lVert X\rVert_{F}.$ (4)
+Let $\Pi_{\perp}$ be the orthogonal projection onto $\mathrm{Null}(\Delta_{N})^{\perp}$ (equivalently, the zero-mean subspace along the sequence dimension). Then
+$\lVert\Pi_{\perp}\mathcal{S}_{\alpha}(X)\rVert_{F}$ $\leq\kappa(\alpha,L)\,\lVert\Pi_{\perp}X\rVert_{F},$ (5)
+$\kappa(\alpha,L)$ $=\max_{k\geq 1}\bigl(1+\alpha\lambda_{k}\bigr)<1.$
+The constraint $\alpha<\tfrac{1}{2}$ is a discrete stability budget analogous to the CFL condition for explicit diffusion: it ensures non-expansiveness and strict attenuation of non-constant (high-frequency) components, while leaving the constant mode unchanged.
+To quantify locality, we use the discrete Dirichlet energy.
+Definition 4.3 (Discrete Dirichlet Energy).
+For $X\in\mathbb{R}^{L\times d}$,
+$\mathcal{E}(X)\triangleq\frac{1}{2}\langle X,-\Delta_{N}X\rangle=\frac{1}{2}\sum_{j=1}^{d}\sum_{i=1}^{L-1}(X_{i+1,j}-X_{i,j})^{2},$ (6)
+which measures local roughness along the sequence axis.
+Theorem 4.4 (Monotone Energy Dissipation).
+For any $0\leq\alpha<\tfrac{1}{2}$ and $X\in\mathbb{R}^{L\times d}$,
+$\mathcal{E}(\mathcal{S}_{\alpha}(X))\leq\mathcal{E}(X),$ (7)
+with strict inequality whenever $\Pi_{\perp}X\neq 0$ and $\alpha>0$.
+Corollary 4.5 (Multi-Scale Stability Under a Global Budget).
+Let $\mathcal{S}_{\boldsymbol{\alpha}}$ be the multi-scale diffusion operator defined above.
+If $\sum_{k=1}^{K}\alpha_{k}<\tfrac{1}{2}$, then $\mathcal{S}_{\boldsymbol{\alpha}}$ is spectrally non-expansive and dissipates $\mathcal{E}(\cdot)$ monotonically.
+Practical takeaway.
+With a strict global budget ($\alpha<\tfrac{1}{2}$ or $\sum_{k}\alpha_{k}<\tfrac{1}{2}$), PDE diffusion is a safe architectural primitive: it cannot amplify representations and it monotonically reduces local roughness.
+4.2 Layer II: Compositional Perturbation Bounds
+We next study how diffusion perturbs a Transformer computation when inserted at different locations.
+Let $f_{\ell}:\mathbb{R}^{L\times d}\to\mathbb{R}^{L\times d}$ denote sub-modules (LayerNorm, attention, MLP), and consider the composition $F=f_{m}\circ\cdots\circ f_{1}$.
+Assumption 4.6 (Local Lipschitz Regularity).
+For each $\ell$, there exists $L_{\ell}>0$ such that in a neighborhood of interest,
+$\lVert f_{\ell}(X)-f_{\ell}(Y)\rVert_{F}\leq L_{\ell}\lVert X-Y\rVert_{F}.$ (8)
+Define the variant with diffusion inserted after $f_{j}$:
+$F^{(j)}\triangleq f_{m}\circ\cdots\circ f_{j+1}\circ\mathcal{S}_{\alpha}\circ f_{j}\circ\cdots\circ f_{1}.$ (9)
+Theorem 4.7 (Perturbation Propagation Bound).
+Under Assumption 4.6, for any input $X$ and $Z_{j}=(f_{j}\circ\cdots\circ f_{1})(X)$,
+$\lVert F^{(j)}(X)-F(X)\rVert_{F}$ $\leq\left(\prod_{\ell=j+1}^{m}L_{\ell}\right)\alpha\,\lVert\Delta_{N}Z_{j}\rVert_{F}$ (10)
+$\leq\left(\prod_{\ell=j+1}^{m}L_{\ell}\right)\alpha\,\sqrt{2\mathcal{E}(Z_{j})}.$
+This bound isolates two insertion-dependent quantities: the local roughness at the insertion point, captured by $\mathcal{E}(Z_{j})$, and the downstream sensitivity, captured by $\prod_{\ell=j+1}^{m}L_{\ell}$.
+In particular, diffusion is guaranteed to be a small perturbation whenever both $\mathcal{E}(Z_{j})$ and the downstream amplification are controlled; conversely, sensitivity to insertion position is expected when either factor is large.
+Practical takeaway.
+The same diffusion operator can help or hurt depending on where it is inserted: its effect scales with insertion-point roughness and downstream amplification, explaining why early insertion is typically most beneficial.
+4.3 Layer III: Diffusion–Attention Non-Commutativity as a Diagnostic
+Finally, we formalize the difference between pre-attention and post-attention diffusion through an exact commutator identity.
+Let $\mathcal{A}$ denote the (possibly input-dependent) self-attention operator.
+Proposition 4.8 (Commutator Identity).
+For any $X$, define
+$Y_{\mathrm{pre}}$ $\triangleq\mathcal{A}(\mathcal{S}_{\alpha}(X)),$ (11)
+$Y_{\mathrm{post}}$ $\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X)).$
+Then
+$Y_{\mathrm{post}}-Y_{\mathrm{pre}}$ $=[\mathcal{S}_{\alpha},\mathcal{A}]X$ (12)
+$\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X))-\mathcal{A}(\mathcal{S}_{\alpha}(X)),$
+and hence $\lVert Y_{\mathrm{post}}-Y_{\mathrm{pre}}\rVert_{F}=\lVert[\mathcal{S}_{\alpha},\mathcal{A}]X\rVert_{F}$.
+To interpret this quantity, we use a local linearization at a fixed input $X$: approximate attention as a mixing operator along the sequence dimension, $\mathcal{A}(X)\approx A(X)X$, where $A(X)\in\mathbb{R}^{L\times L}$ is row-stochastic.
+Substituting into the commutator yields the heuristic form
+$\mathcal{S}_{\alpha},\mathcal{A}]X\approx\alpha\bigl(\Delta_{N}A(X)-A(X)\Delta_{N}\bigr)X,$ (13)
+which makes explicit that non-commutativity arises from a mismatch between local geometry ($\Delta_{N}$) and non-local routing ($A(X)$).
+We emphasize that Eq.
+(13) is a diagnostic approximation rather than a global guarantee: self-attention is nonlinear and input-dependent, so this layer provides interpretable guidance, not a universal bound.
+Practical takeaway.
+Large diffusion–attention mismatch (a large commutator) indicates structural conflict between local smoothing and non-local routing, and motivates avoiding post-attention diffusion when attention induces strong non-local mixing.
+4.4 Design Implications
+The analysis suggests three practical implications.
+First, PDE diffusion provides a certified local geometric prior (Layer I) that is stable under a strict budget $\alpha<\tfrac{1}{2}$.
+Second, the effect of inserting PDE diffusion is controlled by insertion-point roughness and downstream amplification (Layer II), giving a principled notion of “position sensitivity” without relying on empirical observations.
+Third, the commutator $\mathcal{S}_{\alpha},\mathcal{A}]$ exactly measures the representational discrepancy between pre- and post-attention diffusion, and its linearized form attributes large discrepancies to incompatible local and non-local mixing patterns (Layer III).
+Together, these statements justify preferring PDE diffusion placements that regularize representations before aggressive non-local routing, while treating post-routing smoothing as potentially disruptive when the induced commutator is large.
 5 Experiments
-This section aims to empirically validate the core claims of our theoretical framework through a series of rigorous analyses.
-Our experimental design follows a clear logical progression: first, we systematically compare different integration strategies to pinpoint the optimal mechanism for the Adaptive PDE Diffusion Layer, thereby adjudicating the theoretical debate from Section 4.
-Second, we conduct an ablation study to verify the effectiveness of the multi-scale dynamics.
-All experiments are conducted on a challenging long-sequence benchmark to provide a stringent test of our theory.
+This section empirically validates the core claims of our operator-theoretic framework.
+We first compare insertion positions to test the placement predictions implied by the compositional and non-commutativity analyses (Section 4).
+We then ablate the multi-scale design under a fixed global stability budget to test the frequency-coverage hypothesis.
+All experiments are conducted on a challenging long-sequence benchmark to provide a stringent test of the theory.
 5.1 Experimental Setup
-We conduct all experiments on the Long Range Arena (LRA) benchmark Tay et al.
-(2020), which includes five tasks spanning diverse modalities and challenges.
-Our experimental models integrate the proposed Adaptive PDE Diffusion Layer into a strong, optimized vanilla Transformer baseline.
-We systematically evaluate seven distinct integration positions, as detailed in Table 5, to determine the optimal placement.
+We conduct all experiments on the Long Range Arena (LRA) benchmark (Tay et al., 2020b, a), which includes five tasks spanning diverse modalities and challenges.
+Our models integrate the proposed PDE diffusion layer into a strong, optimized vanilla Transformer baseline, following common architectural and training practices (Devlin et al., 2018; Liu et al., 2019; Raffel et al., 2019).
+We systematically evaluate the seven canonical insertion positions defined in Table 1 to determine where locality should be introduced.
 All models were trained on NVIDIA A100-80GB GPUs, with reproducibility ensured by fixing random seeds.
-Comprehensive details regarding the datasets, task-specific model configurations, unified training hyperparameters, and the software/hardware environment are provided in Appendix A to ensure full transparency and reproducibility of our work.
-Table 5: Design space of PDE integration positions in Transformer architectures.
-Pos Integration Point Description
-1 After Embedding Applied right after the input embedding layer, before any Transformer block, so diffusion acts on raw semantic representations.
-2 After MLP Inserted after each block’s MLP sub-layer; diffusion runs on features that have passed through attention and feed-forward transformations.
-3 Layer Diffusion Placed between consecutive Transformer layers to promote inter-layer information flow across depth.
-4 Before LayerNorm Applied just before LayerNorm in every sub-layer, operating on unnormalized features.
-5 In Attention Integrated into the attention mechanism itself—e.g., on attention weights or intermediate values—during computation.
-6 Head Diffusion Acts across attention heads within the same layer, enabling information exchange among heads.
-7 After Attention Inserted immediately after the self-attention sub-layer and before the MLP, diffusing attention-processed representations.
-4.2 Main Results: Pinpointing the Optimal PDE Mechanism
+Additional dataset, configuration, and training details are provided in Appendix A.
+5.2 Validating Position-Dependent Predictions
+Before presenting the results, we clarify how our experiments are designed to directly test the three-layer operator-theoretic framework introduced in Section 4.
+We reuse the canonical diffusion insertion strategies defined in Table 1, spanning early (After-Embedding) to late (After-Attention) placements along with several intermediate variants.
+Tested predictions.
+Our analysis yields three simple, falsifiable placement predictions.
+(Pred-1) Early PDE diffusion (e.g., After-Embedding) should be most beneficial, since representations are rough and regularization has maximal leverage (Layer II).
+(Pred-2) Post-attention PDE diffusion should be harmful due to a structural double-mixing conflict between local smoothing and non-local routing (Layer III).
+(Pred-3) Intermediate placements should yield smaller gains as representations become progressively smoother through depth (Layer II).
+Main results.
+Across five Long Range Arena tasks, the results strongly validate the position-dependent predictions of our framework.
+Consistent with Pred-1 (Layer II), early PDE diffusion yields the best performance: inserting PDE diffusion immediately after the embedding layer (P1) improves average accuracy from $58.62\%$ to $62.69\%$ (+4.1+4.1 points).
+Consistent with Pred-2 (Layer III), post-attention PDE diffusion is harmful: After-Attention insertion (P7) drops to $56.17\%$ (2.5-2.5 points), matching the predicted double-mixing conflict between local smoothing and non-local routing.
+Intermediate insertion strategies (P2–P4) produce smaller gains, consistent with Pred-3 (Layer II) and the diminishing-leverage interpretation as representations become smoother through depth.
+Overall, insertion position is not a tunable detail but a theoretically constrained design choice governed by local–global operator interaction.
 Table 2: Average accuracy on the LRA benchmark for different PDE integration positions. Detailed per-task scores are in Appendix B.
@@
 PDE-After-Attention 0.5617
-To systematically adjudicate the theoretical debate from Section 4 regarding the optimal integration position, we conducted a comprehensive comparison of seven different PDE configurations.Table 6 summarizes the average performance of each configuration on the LRA benchmark.
-The results clearly show that the ‘After Embedding‘ position achieves a compellingly superior performance, outperforming the strong baseline by 4.07 percentage points.
-This finding provides powerful empirical support for our proactive pre-processing hypothesis.
-It reveals a core mechanism: applying Pillar II (Diffusion Geometry) directly on the initial, potentially noisy semantic manifold enforces a local smoothness inductive bias that provides a more regularized and robust foundation for all subsequent non-local interactions (i.e., self-attention).
-This strategy of “semantic regularization” at the source is substantially more effective than “reactive refinement” later in the information flow (e.g., ‘After MLP‘).
-Concurrently, the failure of the ‘After Attention‘ position is equally insightful.
-It delineates the boundaries of our method: once a component (self-attention) has learned valuable sparse association patterns, forcibly imposing dense local smoothing can act as a destructive interference.
-This observation reinforces the functional complementarity between PDE diffusion and self-attention discussed in Section 4.1—they are orthogonal mechanisms, and their simple serial composition is not always optimal, pointing towards future explorations of more sophisticated fusion strategies as suggested by Pillar IV.
 Figure 1: Overall performance comparison on the LRA benchmark. Error bars show standard deviation across five runs (n=5n{=}5).
-4.3 Ablation Study: Validating Multi-Scale Dynamics
-Having identified ‘After Embedding‘ as a superior integration strategy, we conducted a further ablation study to validate the efficacy of Pillar III (Multi-Scale Dynamics).
-We selected the top three performing positions and compared the effects of different diffusion scales ($h=1,2,4$) versus the adaptive multi-scale combination on the ListOps task.
-The results are presented in Table 7.
-Table 7: Ablation study results for the multi-scale diffusion mechanism on the ListOps task.
-Position Fast Medium Slow Multi-
-(hh=1) (hh=2) (hh=4) scale
-After Emb. 0.3960 0.3940 0.3990 0.4080
-After MLP 0.3900 0.3930 0.3910 0.4010
-Layer Diff. 0.3850 0.3890 0.3910 0.3970
-The results are remarkably consistent: for all tested positions, the adaptive ‘Multi-scale‘ configuration outperforms any single scale.
-This strongly corroborates our theoretical hypothesis that real-world sequential data contains structural information at varying granularities.
-By enabling the model to adaptively combine different diffusion processes—capturing high-frequency local details via short-range diffusion and establishing low-frequency global trends via long-range diffusion—it can learn more comprehensive and robust feature representations.
-This empirical success is also in strong agreement with frequency analysis in signal processing theory (see Appendix C.3).
-Figure 4: Detailed results of the multi-scale PDE ablation study on the ListOps task, broken down by PDE position.
+5.3 Validating Multi-Scale Frequency Coverage
+We next evaluate the prediction of Corollary 4.5, which states that multi-scale PDE diffusion preserves intrinsic stability under the global budget constraint $\sum_{k}\alpha_{k}<1/2$ while providing broader frequency coverage than any single scale.
+To test this claim, we focus on the ListOps task and compare three single-scale variants (Fast h=1h{=}1, Medium h=2h{=}2, Slow h=4h{=}4) against an adaptive multi-scale configuration combining h=1,2,4h{=}1,2,4, evaluated at the top three insertion positions.
+All settings enforce the same total budget $\sum_{k}\alpha_{k}=0.48<1/2$ to isolate the effect of frequency coverage.
+As shown in Table 7 (Appendix), adaptive multi-scale diffusion consistently outperforms the best single-scale baseline across all positions, yielding improvements of +0.90+0.90 percentage points for After-Embedding, +0.80+0.80 for After-MLP, and +0.60+0.60 for Layer-Diffusion.
+These gains support the theoretical prediction that different diffusion scales capture complementary frequency bands (see Appendix C.3), and that their combination provides more uniform spectral coverage without violating stability guarantees.
+The consistency of improvements across insertion positions suggests that frequency heterogeneity is a general characteristic of Long Range Arena tasks, reinforcing multi-scale diffusion as a robust design principle rather than a position-specific optimization.
+Figure 2: Multi-scale ablation on ListOps. The adaptive multi-scale configuration consistently outperforms all single-scale variants, validating Corollary 4.5 under a fixed global stability budget.
 5.4 Conclusion of Experiments
-In summary, our empirical investigation systematically validates the core tenets of our theoretical framework, culminating in a clear and powerful conclusion: **introducing a local smoothness inductive bias to Transformers is an effective design principle, and the optimal strategy to achieve this is through proactive pre-processing of input representations with an adaptive, multi-scale PDE diffusion layer.** This conclusion not only demonstrates the effectiveness of the proposed PDE-Transformer but, more importantly, it illuminates the profound complementary relationship between the global information aggregation of self-attention and the local structural smoothing of PDE diffusion.
-Our work offers a new, principled, and rigorously validated approach for future sequence model design: ensure a robust understanding of underlying local structure before modeling complex long-range dependencies.
+Our experiments on Long Range Arena (Tay et al., 2020b, a) provide a controlled validation of the proposed operator-theoretic framework across all three analytical layers.
+First, consistent and stable training under $\alpha=0.1<1/2$ confirms that the spectral budget from Layer I (intrinsic stability) is practically viable.
+Second, position-dependent comparisons (Table 5 in the appendix) align closely with the compositional bound in Theorem 4.7: early-stage insertion (P1) yields a +4.1+4.1 percentage point improvement, while late-stage insertion after attention (P7) results in a 2.5-2.5 point degradation, supporting the predicted roughness–sensitivity trade-off from Layer II.
+Third, the systematic degradation observed at P7 is consistent with the Layer III non-commutativity analysis, lending empirical support to the double-mixing hypothesis implied by the commutator $\mathcal{S},\mathcal{A}]$, although we do not directly measure commutator norms in this work.
+Finally, the multi-scale extension predicted by Corollary 4.5 is supported by consistent gains in the ablation study (Table 7 in the appendix).
+At the same time, this empirical validation is intentionally scoped: LRA is a controlled benchmark rather than a large-scale language modeling setting (Brown et al., 2020); we treat $\mathcal{E}(Z_{j})$ and $\mathcal{S},\mathcal{A}]$ as explanatory quantities rather than directly measured causal signals; and multi-scale ablations are limited to a single task (ListOps).
+Overall, the evidence supports the main claim that operator-theoretic reasoning can predict insertion-position sensitivity, while leaving broader scaling and mechanistic measurement as open directions.
 6 Conclusion
-PDE-Transformer introduces a fundamentally new way to think about sequence modeling by reframing the Transformer’s forward pass as the discretization of a continuous reaction–diffusion system; by deriving an energy functional whose gradient flow yields a PDE with four natural components—non-local interaction for self-attention, local reaction for feed-forward networks, diffusion for positional smoothing, and stability control for normalization—it provides a unified theoretical lens that explains why residual connections and layer normalization are not mere engineering tricks but necessary mechanisms for well-posedness and stability, bridging gaps in our understanding of long-range dependency modeling and revealing that diffusion processes yield polynomial-decay kernels that more effectively capture distant interactions than the exponential decay inherent in standard attention; building on this theory, an Adaptive PDE Diffusion Layer is designed that approximates the Laplacian via second-order finite differences with a learnable diffusion coefficient to adaptively enforce local smoothness, and when integrated into the Transformer at the optimal location—immediately after the embedding layer—this lightweight module achieves a 4.1 pp average accuracy boost on the Long Range Arena benchmark, with a multi-scale version delivering further gains, while a thorough theoretical analysis of seven integration strategies and extensive experiments demonstrate that continuous PDE smoothing and global self-attention are highly complementary, offering a principled, efficient route to robust, long-sequence modeling.
-6 Ethics Statement
+We introduced an operator-theoretic framework for understanding where to add locality in Transformers.
+Under a simple stability budget ($\alpha<1/2$), a PDE diffusion layer provides a certified non-expansive local prior; when composed with the Transformer backbone, its impact becomes position-dependent via a roughness–sensitivity trade-off; and when composed with attention, non-commutativity explains why post-attention smoothing can be structurally harmful.
+On Long Range Arena (Tay et al., 2020b, a), these predictions are borne out: inserting PDE diffusion after embedding improves average accuracy by +4.1+4.1 points, while inserting PDE diffusion after attention degrades performance by 2.5-2.5 points, and a multi-scale PDE variant yields consistent gains under the same global budget.
+More broadly, the main contribution is a transferable template for reasoning about local–global compositions, separating provable guarantees from compositional bounds and diagnostic heuristics.
+Scope and limitations.
+Our claims are scoped to (i) a spectrally-budgeted PDE diffusion operator along the sequence dimension and (ii) controlled long-sequence benchmarks (LRA) with moderate model sizes and contexts.
+We do not directly measure geometric diagnostics (e.g., commutator norms) in the current study, and the multi-scale ablation is limited in task coverage.
+Future work.
+Promising directions include mechanistic validation by measuring roughness and commutator statistics across depth; scaling to language modeling benchmarks and larger models; integrating PDE diffusion more tightly inside attention computations; and applying the same operator-theoretic lens to other hybrid local–global modules (e.g., convolution–attention and state-space–attention).
+7 Impact Statements
 Our research does not involve human subjects, personally identifiable information, or sensitive data.
@@
 We acknowledge that any potential societal impact of large-scale models, such as fairness, bias, or privacy concerns, lies beyond the specific scope of this paper but remains an important consideration for future research.
-7 Reproducibility Statement
-We place strong emphasis on the reproducibility of our results.
-In the main text, we clearly describe our proposed theoretical framework, its derivation, and the algorithmic implementation details.
-A complete set of proofs for theoretical results is provided in the Appendix.
-Hyperparameter configurations, training setups, and ablation study details are explicitly reported in the Experiments section.
-All datasets used are standard and publicly available, and their preprocessing steps are fully documented in the supplementary material.
-To facilitate replication, we will release anonymized source code and scripts for reproducing all experiments upon publication.
-8 LLM Usage Statement
-In accordance with ICLR guidelines on the disclosure of Large Language Model (LLM) usage, we clarify that no LLM contributed substantively to the conception, methodology, or analysis presented in this paper.
-LLMs (e.g., ChatGPT) were used exclusively as auxiliary tools for writing assistance, language refinement, and stylistic editing.
-All technical content, theoretical contributions, experimental design, and analysis were conceived, implemented, and validated entirely by the authors.
-The role of LLMs was limited to improving clarity of presentation and does not rise to the level of authorship or contribution under ICLR policy.
-Appendix A Appendix: Rigorous Mathematical Proofs
-This appendix provides the rigorous mathematical proofs for the theoretical frameworks underpinning the PDE-Transformer.
-These proofs establish the model’s stability, long-range modeling capability, multi-scale representation power, and component coordination.
-A.1 A. Theoretical Framework I: Lyapunov Stability
-This framework analyzes the stability of the model’s underlying gradient flow system using Lyapunov’s second method.
-Theorem A.1 (Energy Monotonicity and Global Stability).
-Consider the energy functional $E[u]$ within the PDE-Transformer, defined as:
-$E[u]=\int_{\Omega}\left(\frac{1}{2}|\nabla u|^{2}+F(u)+\frac{\lambda}{2}|u-u_{0}|^{2}\right)dx,$
-which corresponds to the gradient flow dynamics:
-$\frac{\partial u}{\partial t}=-\frac{\delta E}{\delta u}=\Delta u-F^{\prime}(u)-\lambda(u-u_{0}).$
-If the potential $F(u)$ is strictly convex, satisfying $F^{\prime\prime}(u)\geq\mu>0$, the system is globally asymptotically stable.
-Proof.
-We establish stability by demonstrating that the energy functional $E[u]$ serves as a strict Lyapunov function for the system.
-Step 1: Construct the Lyapunov Function.
-We select the Lyapunov function $V[u]=E[u]$.
-By choosing an appropriate reference, we can assume $F(u)\geq 0$.
-Since $\lambda>0$, the functional $V[u]$ is bounded from below by 0.
-Step 2: Compute the Time Derivative. We compute the time derivative of $V[u]$ along the system’s trajectories:
-$\frac{dV}{dt}$ $=\int_{\Omega}\frac{\delta E}{\delta u}\frac{\partial u}{\partial t}\,dx$
-$=\int_{\Omega}\frac{\delta E}{\delta u}\left(-\frac{\delta E}{\delta u}\right)\,dx$
-$=-\int_{\Omega}\left\|\frac{\delta E}{\delta u}\right\|^{2}\,dx\leq 0.$
-Equality holds if and only if $\frac{\delta E}{\delta u}=0$, which defines the equilibrium state of the system.
-Step 3: Apply LaSalle’s Invariance Principle.
-The condition $F^{\prime\prime}(u)\geq\mu>0$ ensures that $E[u]$ is strictly convex, guaranteeing a unique global minimum.
-Since $V[u]$ is bounded below and its time derivative is negative semi-definite, all trajectories must converge to the largest invariant set where $\frac{dV}{dt}=0$.
-This set consists solely of the unique global minimum.
-Thus, the system is globally asymptotically stable.
-∎
-Theorem A.2 (Exponential Decay of the Gradient Norm).
-Under the conditions of Theorem A.1, the $L^{2}$-norm of the gradient decays at an exponential rate:
-$\left\|\frac{\delta E}{\delta u}(t)\right\|_{L^{2}}\leq\left\|\frac{\delta E}{\delta u}(0)\right\|_{L^{2}}e^{-\mu t}.$
-Proof.
-Step 1: Define the Evolution Equation for the Gradient. Let $v=\frac{\partial u}{\partial t}=-\frac{\delta E}{\delta u}$. Taking the time derivative of $v$ yields:
-$\frac{\partial v}{\partial t}=\frac{\partial}{\partial t}\left(-\frac{\delta E}{\delta u}\right)=\Delta v-F^{\prime\prime}(u)v-\lambda v.$
-Step 2: Analyze the Dynamics of the Gradient Norm. We examine the time derivative of the squared $L^{2}$-norm of $v$:
-$\frac{1}{2}\frac{d}{dt}\|v\|_{L^{2}}^{2}$ $=\int_{\Omega}v\frac{\partial v}{\partial t}\,dx$
-$=\int_{\Omega}v(\Delta v-F^{\prime\prime}(u)v-\lambda v)\,dx$
-$=-\|\nabla v\|_{L^{2}}^{2}-\int_{\Omega}F^{\prime\prime}(u)v^{2}\,dx-\lambda\|v\|_{L^{2}}^{2}.$
-Step 3: Establish the Differential Inequality. Using the condition $F^{\prime\prime}(u)\geq\mu$ and dropping the non-positive term $\|\nabla v\|_{L^{2}}^{2}$, we obtain:
-$\frac{d}{dt}\|v\|_{L^{2}}^{2}\leq-2\mu\|v\|_{L^{2}}^{2}.$
-Applying Grönwall’s inequality to this differential inequality yields $\|v(t)\|_{L^{2}}^{2}\leq\|v(0)\|_{L^{2}}^{2}e^{-2\mu t}$. Taking the square root of both sides completes the proof. ∎
-A.2 B. Framework II: Diffusion Geometry
-Theorem A.3 (Polynomial Decay of the Heat Kernel).
-The heat kernel $K_{t}(x,y)$ corresponding to the 1D discrete Laplacian operator exhibits an asymptotic behavior described by a Gaussian:
-$|K_{t}(x,y)|\leq Ct^{-1/2}\exp\left(-\frac{|x-y|^{2}}{4Dt}\right).$
-This implies a fundamentally slower, polynomial-like decay of influence over long distances compared to the exponential decay of standard self-attention mechanisms.
-Proof.
-Step 1: Spectral Representation. The proof relies on the spectral representation of the heat kernel:
-$K_{t}(x,y)=\sum_{k}e^{-\lambda_{k}t}\phi_{k}(x)\phi_{k}(y),$
-where $\lambda_{k}$ and $\phi_{k}$ are the eigenvalues and eigenfunctions of the discrete Laplacian.
-Step 2: Asymptotic Analysis.
-For large time $t$ or small grid spacing, the discrete sum can be approximated by an integral via the Poisson summation formula.
-This integral evaluates to the continuous heat kernel, which is a Gaussian function.
-While the Gaussian itself has an exponential tail, its integrated effect over time results in a polynomial decay of influence, which is fundamentally slower than the direct exponential decay $e^{-c|x-y|}$ found in typical attention mechanisms.
-∎
-A.3 C. Framework III: Multi-Scale Dynamics
-Theorem A.4 (Multi-Scale Approximation Error Bound).
-Let $\Delta_{\text{multi}}=\sum_{k=1}^{K}\alpha_{k}\Delta_{h_{k}}$ be a multi-scale approximation of the ideal Laplacian $\Delta$.
-By selecting scales $\{h_{k}\}$ geometrically and optimizing weights $\{\alpha_{k}\}$, the approximation error is bounded by:
-$\left\|\Delta_{\text{multi}}-\Delta\right\|_{\mathcal{L}(H^{s},H^{s-2})}\leq CK^{-p},$
-where $p\geq 1$ is the order of convergence.
-Proof.
-This theorem is a result from numerical analysis and approximation theory.
-The problem can be framed as approximating the function $f(\omega)=-\omega^{2}$ (the symbol of the ideal Laplacian) with a linear combination of basis functions $\hat{\Delta}_{h_{k}}(\omega)=-\frac{4}{h_{k}^{2}}\sin^{2}(\frac{\omega h_{k}}{2})$.
-By choosing scales $h_{k}$ in a geometric progression, the basis functions effectively tile the frequency domain.
-Techniques from Chebyshev approximation or spline theory can then be used to show that the $L^{2}$ error of the best approximation decreases polynomially with the number of scales $K$.
-∎
-A.4 D. Framework IV: Multi-Component Coupling
-Theorem A.5 (Synchronization of Coupled Systems).
-Consider a system of $H$ coupled PDEs for multiple attention heads:
-$\frac{\partial u_{i}}{\partial t}=\alpha_{i}\Delta u_{i}+\sum_{j\neq i}\beta_{ij}(u_{j}-u_{i})+f_{i}(u_{i}).$
-If the coupling graph defined by $\beta_{ij}>0$ is connected, the system asymptotically synchronizes, i.e., $\lim_{t\to\infty}\|u_{i}(t)-u_{j}(t)\|=0$ for all pairs $(i,j)$.
-Proof.
-Step 1: Construct the Lyapunov Function. We construct a Lyapunov function representing the total disagreement in the system:
-$V=\frac{1}{2}\sum_{i,j}\|u_{i}-u_{j}\|_{L^{2}}^{2}.$
-This can be expressed compactly using the graph Laplacian $L$ of the coupling network as $V=\sum_{i}\int u_{i}^{T}Lu_{i}\,dx$.
-Step 2: Analyze the Time Derivative.
-The time derivative $\frac{dV}{dt}$ can be shown to be negative semi-definite.
-The coupling term contributes a term proportional to $\lambda_{2}(L)V$, where $\lambda_{2}(L)$ is the algebraic connectivity (the second smallest eigenvalue) of the graph Laplacian $L$.
-Step 3: Conclude Convergence.
-If the nonlinear terms $f_{i}$ are Lipschitz continuous, we can establish that $\frac{dV}{dt}\leq-\mu V$ for some $\mu>0$ that depends on $\lambda_{2}(L)$.
-By Grönwall’s inequality, $V(t)$ decays to zero exponentially, proving synchronization.
-∎
-Appendix B Appendix
-This appendix provides supplementary material to support the main paper.
-Section A details the complete experimental setup, ensuring full reproducibility.
-Section B presents the unabridged experimental data and further analysis.
-Section C contains additional supplementary analyses, including computational costs and robustness checks.
-B.1 Detailed Experimental Setup
-B.1.1 Datasets and Tasks.
-We evaluate our PDE-enhanced Transformer models on the Long Range Arena (LRA) benchmark Tay et al.
-(2020), which consists of five challenging tasks designed to assess the ability of sequence models to capture long-range dependencies:
+Appendix A Supplementary Material
+This appendix provides supplementary material to support the main paper, including experimental setup details, extended results, and additional analyses.
+A.1 Detailed Experimental Setup
+A.1.1 Datasets and Tasks.
+We evaluate our PDE-enhanced Transformer models on the Long Range Arena (LRA) benchmark (Tay et al., 2020b, a), which consists of five challenging tasks designed to assess the ability of sequence models to capture long-range dependencies:
   • ListOps: A synthetic task requiring hierarchical reasoning over sequences of up to 2,000 tokens, with 10 classes.
@@
   • Image Classification: CIFAR-10 classification on $\times32$ images, flattened to 1,024-token sequences, with 10 classes.
-B.1.2 Software and Hardware Environment.
+A.1.2 Software and Hardware Environment.
 All experiments were conducted on a high-performance computing cluster equipped with NVIDIA A100-80GB GPUs.
@@
 Reproducibility was ensured by fixing the random seeds for Python (42), NumPy (0), and PyTorch (1) across all runs.
-B.1.3 Model Hyperparameters.
+A.1.3 Model Hyperparameters.
 All experiments use Transformer configurations optimized for A100-80GB GPUs.
@@
 MLP Hidden Dim 512 (4x model dimension)
-B.1.4 The detailed PDE operation
+A.1.4 The detailed PDE operation
 Each PDE operation implements the discrete diffusion equation:
 $\frac{\partial u}{\partial t}=\alpha\nabla^{2}u$ (14)
-where $\alpha$ is a learnable diffusion coefficient initialized to 0.1, and $\nabla^{2}$ is approximated using the finite difference method with reflective boundary conditions:
+where $\alpha$ is a learnable diffusion coefficient initialized to 0.1, and $\nabla^{2}$ is approximated using the finite difference method with replicate (Neumann) boundary conditions:
 $\nabla^{2}u_{i}\approx u_{i+1}-2u_{i}+u_{i-1}$ (15)
@@
 $u^{(t+1)}=\text{LayerNorm}(u^{(t)}+\alpha\nabla^{2}u^{(t)})$ (16)
-B.2 Complete Experimental Data and Analysis
+A.2 Complete Experimental Data and Analysis
 This section provides the unabridged data from our experiments, offering a more granular view of the results presented in the main paper and a deeper analysis of performance trends.
-B.2.1 Main Experiment: PDE Integration Positions.
+A.2.1 Main Experiment: PDE Integration Positions.
 Figure 3: Seven PDE integration positions in Transformer architecture.
@@
 To better visualize these results, Figure 4 shows the per-task bar charts, while Figure 5 offers a heatmap for quick comparative analysis.
-These detailed results highlight task-specific sensitivities; for instance, the superiority of ‘PDE-After-Embed‘ is particularly pronounced on the PathFinder and Text Classification tasks, which heavily rely on spatial and semantic reasoning, respectively.
+These detailed results highlight task-specific sensitivities; for instance, the superiority of PDE-After-Embed is particularly pronounced on the PathFinder and Text Classification tasks, which heavily rely on spatial and semantic reasoning, respectively.
 Table 5: Detailed performance comparison of different PDE integration positions on the LRA benchmark. Accuracy scores are reported for each task. Best performance per task is highlighted in bold.
@@
 Figure 5: Performance heatmap of all model configurations across the five LRA benchmark tasks. Warmer colors indicate higher performance.
-B.2.2 Task-wise Improvement Analysis.
+A.2.2 Task-wise Improvement Analysis.
 Table 6 further quantifies the differences observed in the previous section by detailing the absolute ($\Delta$) and relative (%) performance gains of each configuration over the baseline for every task.
@@
 After Attention +0.0000 +0.00 -0.0038 -0.59 -0.0001 -0.01 -0.1336 -19.04 +0.0148 +3.74
-B.2.3 Ablation Study: Multi-Scale Dynamics.
+A.2.3 Ablation Study: Multi-Scale Dynamics.
 This section provides the complete data for the multi-scale ablation study on the ListOps task.
@@
 Figure 8: Detailed results of the multi-scale PDE ablation study on the ListOps task, broken down by PDE position.
-B.2.4 Cross-Task Performance Consistency Analysis.
+A.2.4 Cross-Task Performance Consistency Analysis.
 To assess the generalizability of our approach, we computed a cross-task performance correlation matrix (Table 8).
@@
 PathFinder $\leftrightarrow$ Image 0.87 0.64 0.69 0.71 +0.12
-B.3 Supplementary Analysis
-B.3.1 Computational Cost Analysis.
+A.3 Supplementary Analysis
+A.3.1 Computational Cost Analysis.
 The PDE Diffusion Layer is computationally efficient.
@@
 Multi-scale (Emb.) +18.4% +3.7%
-B.3.2 Robustness and Sensitivity Analysis.
+A.3.2 Robustness and Sensitivity Analysis.
 To ensure our results were not coincidental, we performed a limited evaluation with multiple random seeds on the top-performing positions for the ListOps task (Table 10).
@@
 Layer Diffusion 0.3850 $\pm$ 0.0035
-B.3.3 Frequency Domain Analysis.
+A.3.3 Frequency Domain Analysis.
 The success of multi-scale diffusion can be understood from a frequency domain perspective.
@@
 Multi-scale 0.64 0.82 1.71 1.80
-B.4 Core Mechanism and Theoretical Analysis
-B.4.1 Theoretical Design Principles.
-This seemingly simple formula embodies two key, principled design choices:
-  1. 1.
-    Update Form Based on Numerical Analysis.
-The residual form $X+\dots$ is not merely a conventional skip connection.
-In numerical analysis, it can be rigorously interpreted as a single-step forward Euler method for solving the diffusion equation $\frac{\partial u}{\partial t}=\alpha_{\text{coeff}}\nabla^{2}u$, where the time step $\Delta t$ is effectively absorbed into the learnable coefficient $\alpha$.
-According to Theorem 1.2 (Exponential Decay of the Gradient Norm), this update form guarantees that the system’s energy gradient converges at an exponential rate, providing a solid mathematical foundation for stability:
-    $\left\|\frac{\delta E}{\delta u}(t+\Delta t)\right\|_{L^{2}}\leq\left\|\frac{\delta E}{\delta u}(t)\right\|_{L^{2}}e^{-\mu\alpha}$ (10)
-    where $\mu$ is a positive definite constant of the system, ensuring the stability of the information propagation process.
-  2. 2.
-    Learnable Adaptive Diffusion Strength.
-The diffusion coefficient $\alpha$ is a learnable scalar parameter, which endows the model with a crucial adaptive capability.
-Based on Theorem 3.1 (Frequency Properties of Multi-Scale Filters), the learning process of $\alpha$ is, in fact, an optimization of the layer’s frequency response function:
-    $\hat{\Delta}_{\text{adaptive}}(\omega)=\alpha\cdot\left(-\frac{4}{h^{2}}\sin^{2}\left(\frac{\omega h}{2}\right)\right)$ (11)
-    The model can autonomously decide the required intensity of information smoothing at each layer based on the task and data, and can even effectively ”turn off” the layer by learning an $\alpha$ close to zero.
-B.4.2 Theoretical Complementarity Analysis.
-The PDE Diffusion Layer introduces a powerful inductive bias of local smoothness into the Transformer model. It forms a profound theoretical complementarity with the self-attention mechanism:
-  • Self-Attention: Performs sparse, content-based global information retrieval.
-  • PDE Diffusion: Performs dense, structure-based local information integration.
-According to Theorem 2.1 (Polynomial Decay of the Heat Kernel), this complementarity is manifested in a fundamental difference in their information propagation patterns.
-For two positions at a distance of |xy||x-y|, their influence decay patterns are:
-  • Attention Weights (Typical): $A(x,y)\sim e^{-c|x-y|}$ (Exponential decay)
-  • Diffusion Kernel: $K_{t}(x,y)\sim|x-y|^{-1}$ (Polynomial decay)
-Polynomial decay provides a stronger long-range connection capability, effectively supplementing the deficiencies of standard attention mechanisms in capturing ultra-long-range dependencies.
-B.4.3 Inherent Positional Awareness.
-The PDE Diffusion Layer provides the model with an inherent, structural sense of position that is orthogonal to traditional positional encodings.
-While standard Transformers rely on externally injected signals (e.g., sinusoidal encodings) to perceive order, our diffusion mechanism, through its intrinsic local stencil, tightly couples the representation of each token to its immediate neighbors, thus naturally encoding the topology and relative order of the sequence.
+Appendix B Detailed Theoretical Proofs
+This appendix provides self-contained proofs of the main stability statements for the PDE diffusion operator used in the paper.
+Throughout, we use Neumann (zero-flux) boundary conditions implemented by
+emphreplication at the endpoints (often called replicate or nearest padding in deep-learning libraries):
+$u_{0}=u_{1},\qquad u_{L+1}=u_{L}.$ (17)
+This convention is consistent with the discrete Neumann Laplacian defined below.
+The reflect padding rule corresponds to a different ghost-point convention and would change the boundary rows; throughout, our analysis and implementation assume Eq.
+(17).
+B.1 Notation and the discrete Neumann Laplacian
+Let $L$ be the sequence length and $d$ the hidden dimension.
+Representations are matrices $X\in\mathbb{R}^{L\times d}$, with channels indexed by $j\in\{1,\dots,d\}$.
+The 1D discrete Neumann Laplacian $\Delta_{N}\in\mathbb{R}^{L\times L}$ acts on a vector $x\in\mathbb{R}^{L}$ by
+$(\Delta_{N}x)_{i}=\begin{cases}x_{2}-x_{1},&i=1,\\ x_{i-1}-2x_{i}+x_{i+1},&2\leq i\leq L-1,\\ x_{L-1}-x_{L},&i=L.\end{cases}$ (18)
+For matrix inputs $X$, $\Delta_{N}$ acts channel-wise along the sequence dimension: $(\Delta_{N}X)_{\cdot j}=\Delta_{N}(X_{\cdot j})$.
+We define the diffusion operator
+$\mathcal{S}_{\alpha}(X)\triangleq(I+\alpha\Delta_{N})X,\qquad 0\leq\alpha<\tfrac{1}{2}.$ (19)
+Unless stated otherwise, $\|\cdot\|_{2}$ denotes the spectral norm and $\|\cdot\|_{F}$ denotes the Frobenius norm.
+B.2 Proof of the spectrum of $\Delta_{N}$
+Lemma B.1 (Spectrum of the discrete Neumann Laplacian).
+$\Delta_{N}$ is real symmetric and negative semi-definite. Its eigenvalues are
+$\lambda_{k}=-4\sin^{2}\!\Bigl(\frac{\pi k}{2L}\Bigr),\qquad k=0,1,\dots,L-1,$ (20)
+so that $\lambda_{k}\in[-4,0]$ and $\lambda_{0}=0$ corresponds to the constant mode.
+Proof.
+(1) Symmetry and semi-definiteness.
+From Eq.
+(18), $\Delta_{N}$ is tri-diagonal with boundary rows $(\Delta_{N}x)_{1}=x_{2}-x_{1}$ and $(\Delta_{N}x)_{L}=x_{L-1}-x_{L}$, hence its matrix form is symmetric.
+Moreover, for any $x\in\mathbb{R}^{L}$,
+[…]
+which implies $\Delta_{N}\preceq 0$.
+(2) Eigenpairs. Consider the cosine family (a DCT-type basis)
+$v^{(k)}_{i}=\cos\!\left(\frac{\pi k}{L}\left(i-\frac{1}{2}\right)\right),\qquad i=1,\dots,L,$ (22)
+for $k=0,1,\dots,L-1$. A direct computation for interior indices uses the identity $\cos(a-b)-2\cos(a)+\cos(a+b)=2(\cos b-1)\cos a$ with $b=\tfrac{\pi k}{L}$ to give
+$(\Delta_{N}v^{(k)})_{i}=2\Bigl(\cos\!\bigl(\tfrac{\pi k}{L}\bigr)-1\Bigr)v^{(k)}_{i}=-4\sin^{2}\!\Bigl(\tfrac{\pi k}{2L}\Bigr)v^{(k)}_{i},\qquad 2\leq i\leq L-1.$ (23)
+The same eigenvalue also satisfies the boundary rows in Eq.
+(18) (consistent with the endpoint replication convention), hence each $v^{(k)}$ is an eigenvector with eigenvalue $\lambda_{k}$ in Eq.
+(20).
+(3) Range. For $k=0$, $\lambda_{0}=0$ and $v^{(0)}$ is constant. For $k\geq 1$, $\sin(\cdot)>0$ implies $\lambda_{k}<0$, and the minimum approaches 4-4. This completes the proof. ∎
+B.3 Proof of spectral non-expansiveness and strict contraction off the nullspace
+Theorem B.2 (Spectral non-expansiveness; strict contraction off the nullspace).
+For $0\leq\alpha<\tfrac{1}{2}$, the diffusion operator $\mathcal{S}_{\alpha}$ in Eq. (19) satisfies
+$\|\mathcal{S}_{\alpha}\|_{2}\leq 1,\qquad\|\mathcal{S}_{\alpha}(X)\|_{F}\leq\|X\|_{F}\ \ \text{for all }X\in\mathbb{R}^{L\times d}.$ (24)
+Let $\Pi_{\perp}$ be the orthogonal projector onto $\mathrm{Null}(\Delta_{N})^{\perp}$ (i.e., the zero-mean subspace along the sequence dimension).
+Then there exists $\kappa(\alpha,L)\in(0,1)$ such that
+$\|\Pi_{\perp}\mathcal{S}_{\alpha}(X)\|_{F}\leq\kappa(\alpha,L)\,\|\Pi_{\perp}X\|_{F}.$ (25)
+Proof.
+Since $\Delta_{N}$ is symmetric (Lemma B.1), it admits an orthogonal eigen-decomposition $\Delta_{N}=U\Lambda U^{\top}$ with $\Lambda=\mathrm{diag}(\lambda_{0},\dots,\lambda_{L-1})$.
+The operator $\mathcal{S}_{\alpha}=I+\alpha\Delta_{N}$ is also symmetric with eigenvalues
+$\mu_{k}=1+\alpha\lambda_{k}.$ (26)
+By Lemma B.1, $\lambda_{k}\in[-4,0]$. For $0\leq\alpha<\tfrac{1}{2}$, we have
+$0\leq 1-4\alpha\leq\mu_{k}\leq 1.$ (27)
+Because $\mathcal{S}_{\alpha}$ is symmetric, its spectral norm equals the maximum absolute eigenvalue: $\|\mathcal{S}_{\alpha}\|_{2}=\max_{k}|\mu_{k}|\leq 1$, proving the first claim.
+Then for any $X$, using the standard inequality $\|MX\|_{F}\leq\|M\|_{2}\|X\|_{F}$ yields $\|\mathcal{S}_{\alpha}(X)\|_{F}\leq\|\mathcal{S}_{\alpha}\|_{2}\|X\|_{F}\leq\|X\|_{F}$.
+For strict contraction off the nullspace, observe that $\mathrm{Null}(\Delta_{N})$ is spanned by the constant eigenvector (mode $k=0$), for which $\mu_{0}=1$.
+On $\mathrm{Null}(\Delta_{N})^{\perp}$ only modes $k\geq 1$ appear, and since $\lambda_{1}=\max_{k\geq 1}\lambda_{k}<0$ we have
+$\kappa(\alpha,L)\triangleq\max_{k\geq 1}\mu_{k}=1+\alpha\lambda_{1}\in(0,1).$ (28)
+Therefore $\|\Pi_{\perp}\mathcal{S}_{\alpha}(X)\|_{F}\leq\kappa(\alpha,L)\|\Pi_{\perp}X\|_{F}$. ∎
+B.4 Proof of monotone Dirichlet-energy dissipation
+Definition B.3 (Discrete Dirichlet energy).
+For $X\in\mathbb{R}^{L\times d}$, define
+$\mathcal{E}(X)\triangleq\tfrac{1}{2}\langle X,-\Delta_{N}X\rangle\;=\;\tfrac{1}{2}\sum_{j=1}^{d}X_{\cdot j}^{\top}(-\Delta_{N})X_{\cdot j}\;=\;\tfrac{1}{2}\sum_{j=1}^{d}\sum_{i=1}^{L-1}(X_{i+1,j}-X_{i,j})^{2}.$ (29)
+Theorem B.4 (Monotone energy dissipation).
+For $0\leq\alpha<\tfrac{1}{2}$ and all $X\in\mathbb{R}^{L\times d}$,
+$\mathcal{E}(\mathcal{S}_{\alpha}(X))\leq\mathcal{E}(X),$ (30)
+with strict inequality whenever $\Pi_{\perp}X\neq 0$ and $\alpha>0$.
+Proof.
+Because $\Delta_{N}$ acts independently across channels, it suffices to prove the claim for a single channel $x\in\mathbb{R}^{L}$ and then sum over $j$.
+Let $\Delta_{N}=U\Lambda U^{\top}$ and write $\hat{x}=U^{\top}x$.
+Then
+$\mathcal{E}(x)=\tfrac{1}{2}\,x^{\top}(-\Delta_{N})x=\tfrac{1}{2}\sum_{k=0}^{L-1}(-\lambda_{k})\,\hat{x}_{k}^{2}.$ (31)
+After diffusion, $\widehat{\mathcal{S}_{\alpha}x}_{k}=(1+\alpha\lambda_{k})\hat{x}_{k}$, hence
+$\mathcal{E}(\mathcal{S}_{\alpha}x)=\tfrac{1}{2}\sum_{k=0}^{L-1}(-\lambda_{k})\,(1+\alpha\lambda_{k})^{2}\,\hat{x}_{k}^{2}.$ (32)
+For $k=0$, $\lambda_{0}=0$ so the summand is zero.
+For $k\geq 1$, $\lambda_{k}>0$ and, by Eq.
+(27), $0\leq(1+\alpha\lambda_{k})^{2}\leq 1$, so each summand in Eq.
+(32) is at most the corresponding summand in Eq.
+(31).
+Summing over $k$ gives (30).
+If $\alpha>0$ and $\Pi_{\perp}x\neq 0$, then there exists some $k\geq 1$ with $\hat{x}_{k}\neq 0$ and $\lambda_{k}<0$, which implies $(1+\alpha\lambda_{k})^{2}<1$ and hence strict inequality. ∎
+B.5 Commutator identity and a linearized attention view
+Proposition B.5 (Commutator identity).
+Let $\mathcal{A}$ be any (possibly nonlinear, input-dependent) attention mapping, and let $\mathcal{S}_{\alpha}$ be the diffusion operator. Define
+$Y_{\mathrm{pre}}\triangleq\mathcal{A}(\mathcal{S}_{\alpha}(X)),\qquad Y_{\mathrm{post}}\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X)).$ (33)
+Then
+$Y_{\mathrm{post}}-Y_{\mathrm{pre}}=[\mathcal{S}_{\alpha},\mathcal{A}]X\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X))-\mathcal{A}(\mathcal{S}_{\alpha}(X)),$ (34)
+and therefore $\|Y_{\mathrm{post}}-Y_{\mathrm{pre}}\|_{F}=\|[\mathcal{S}_{\alpha},\mathcal{A}]X\|_{F}$.
+Proof.
+The identity is purely algebraic by direct expansion of the definitions in Eq. (33). ∎
+Linearized attention heuristic.
+At a fixed input $X$, one may locally approximate attention by a mixing matrix $A(X)\in\mathbb{R}^{L\times L}$ acting along the sequence dimension (e.g., row-stochastic weights at $X$): $\mathcal{A}(X)\approx A(X)X$.
+Under this heuristic,
+$\mathcal{S}_{\alpha},\mathcal{A}]X\approx\alpha\bigl(\Delta_{N}A(X)-A(X)\Delta_{N}\bigr)X.$ (35)
+This approximation is intended as an interpretive diagnostic (not a global linear model of attention).
+B.6 Lipschitz propagation bound and roughness control
+Theorem B.6 (Lipschitz propagation bound).
+Let $F=f_{m}\circ\cdots\circ f_{1}$ be a composition of mappings on $\mathbb{R}^{L\times d}$, and define the variant with diffusion inserted after $f_{j}$:
+$F^{(j)}=f_{m}\circ\cdots\circ f_{j+1}\circ\mathcal{S}_{\alpha}\circ f_{j}\circ\cdots\circ f_{1}.$
+Assume the tail maps $f_{j+1},\dots,f_{m}$ are locally Lipschitz in $\|\cdot\|_{F}$ with constants $L_{j+1},\dots,L_{m}$. For an input $X$, let $Z_{j}=(f_{j}\circ\cdots\circ f_{1})(X)$. Then
+$\|F^{(j)}(X)-F(X)\|_{F}\leq\Bigl(\prod_{\ell=j+1}^{m}L_{\ell}\Bigr)\,\|\bigl(\mathcal{S}_{\alpha}-I\bigr)(Z_{j})\|_{F}.$ (36)
+Moreover,
+$\|\bigl(\mathcal{S}_{\alpha}-I\bigr)(Z_{j})\|_{F}=\alpha\|\Delta_{N}Z_{j}\|_{F}\leq\alpha\sqrt{2\,\mathcal{E}(Z_{j})}.$ (37)
+Proof.
+Define the tail composition $T\triangleq f_{m}\circ\cdots\circ f_{j+1}$. By assumption, $T$ is locally Lipschitz in $\|\cdot\|_{F}$ with constant $\prod_{\ell=j+1}^{m}L_{\ell}$. Therefore,
+$\|F^{(j)}(X)-F(X)\|_{F}=\|T(\mathcal{S}_{\alpha}(Z_{j}))-T(Z_{j})\|_{F}\leq\Bigl(\prod_{\ell=j+1}^{m}L_{\ell}\Bigr)\,\|\mathcal{S}_{\alpha}(Z_{j})-Z_{j}\|_{F},$ (38)
+which is Eq. (36).
+From Eq.
+(37), we have $\mathcal{S}_{\alpha}(Z_{j})-Z_{j}=\alpha\Delta_{N}Z_{j}$.
+It remains to relate $\|\Delta_{N}Z_{j}\|_{F}$ to $\mathcal{E}(Z_{j})$.
+For a single channel vector $z\in\mathbb{R}^{L}$, Lemma B.1 implies
+$\|\Delta_{N}z\|_{2}^{2}=\sum_{k}\lambda_{k}^{2}\,\hat{z}_{k}^{2}\leq 4\sum_{k}(-\lambda_{k})\,\hat{z}_{k}^{2}=8\,\mathcal{E}(z),$ (39)
+because $\lambda_{k}^{2}\leq 4(-\lambda_{k})$ for all $\lambda_{k}\in[-4,0]$.
+Summing this bound over channels yields $\|\Delta_{N}Z_{j}\|_{F}^{2}\leq 4\,\langle Z_{j},-\Delta_{N}Z_{j}\rangle=8\,\mathcal{E}(Z_{j})$, i.e., $\|\Delta_{N}Z_{j}\|_{F}\leq\sqrt{2\,\mathcal{E}(Z_{j})}$.
+∎